### v1

# Quantitative propagation of chaos for 2D stochastic vortex model on the whole space under moderate interactions

Abstract
We derive the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in the sense of relative entropy. The main novelty is the application of the Donsker–Varadhan inequality in the context of moderately interacting particles to handle the nonlinearity, as well as the use of localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms. Moreover, we prove the existence of a suitable solution to the aforementioned model.
MSC2010 subject classification: 49N90, 60H30, 60K35
1 Introduction
In this paper, we establish a quantitative derivation of the stochastic 2D vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
$\mathrm{d}\rho_{t}=\Delta\rho_{t}\,dt+\frac{1}{2}D^{2}\rho_{t}(\sigma\sigma^{\top})_{t}\,dt-\nabla\cdot(\rho_{t}(K\ast\rho_{t}))\,dt-\nabla\rho_{t}\cdot\sigma_{t}\,dB_{t},$ (1)
from a stochastic moderately interacting particle system given by
$dX_{t}^{i,N}$ $=\frac{1}{N}\sum_{k=1}^{N}\left(K\ast V^{N}\right)\left(X_{t}^{i,N}-X_{t}^{k,N}\right)\,dt+\sqrt{2}\,dW_{t}^{i,N}+\sigma_{t}\,dB_{t},\quad$ (2)
where $W_{t}^{i,N}$ and $B_{t}$ are independent standard $\mathbb{R}^{d}$-valued Brownian motions, defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$, $K$ is a singular kernel and $V^{N}$ a suitable scalling.
To the best of the author’s knowledge, this is the first time that (1) is derived for 2D stochastic vortex model, by moderately interacting particles, in the whole space - extending what was known before for bounded kernels in [32], on $\mathbb{R}^{d}$, and in [39], for sub-Coulombian kernels, under periodic boundary conditions - with pathwise quantitative bounds, in the sense of the relative entropy functional.
In the particle system (2), the stochastic term $\sigma_{t}dB_{t}$ accounts for a source of randomness that affects all particles, modeling a shared environmental fluctuation. Models of this nature appear in several contexts, such as problems in statistical mechanics, population dynamics and biology, as well as in the analysis of numerical Monte–Carlo schemes and related applications; see, e.g., [4], [5], [24], [19], and [29].
The main goal of this paper is to investigate the asymptotics in $N$ of the empirical measure defined by
$S_{t}^{N}\doteq\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},\qquad t\geqslant 0,$ (3)
where $\delta_{a}$ is the delta Dirac measure concentrated at $a\in\mathbb{R}^{d}$. For that purpose, considering $V^{N}\doteq N^{\beta}V(N^{\frac{\beta}{d}}\cdot)$, we introduce the mollified empirical measure
$\rho_{t}^{N}\doteq V^{N}\ast S_{t}^{N}=\int_{\mathbb{R}^{d}}V^{N}(\cdot-y)\,S_{t}^{N}(dy),$
in order to apply the relative entropy functional directly to the densities $\rho^{N}$ and $\rho$, since by the Csiszar-Kullback-Pinsker inequality, we get
$\|S^{N}_{t}-\rho_{t}\|_{0}^{2}\lesssim N^{-\frac{\beta}{d}}+\|\rho^{N}_{t}-\rho_{t}\|_{1}^{2}\overset{\text{CKP}}{\lesssim}N^{-\frac{\beta}{d}}+\mathcal{H}\left(\rho^{N}_{t}|\rho_{t}\right),$ (4)
where the metric in the left hand side of (4), is the Kantorovich-Rubinstein metric between two probatility measures. It follows that, if there exists a constant $C=C_{\omega}>0$ such that for any $N\in\mathbb{N}$,
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})\leq C(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}),$ (5)
where $\theta$ is an explicit positive parameter, by (4) and (5), we have, in particular, propagation of chaos for the marginals of the empirical measure of the particle system, see [1, Section 8.3] and [45].
In order to derive (5), we apply for the first time in the context of moderately interacting particles, the Donsker–Varadhan inequality together with the dissipation provided by the Fisher information, to handle the nonlinearity. In addition, we employ localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms in the whole space, as the estimate established in [39], under periodic boundary conditions, does not apply in the present framework. Moreover, we prove the existence of a suitable solution to the equation (1).
Related works
Quantitative estimates for singular interactions have been substantially developed in recent years. The relative entropy method for establishing quantitative propagation of chaos in McKean–Vlasov systems with singular interaction kernels was first introduced in [23]. It was developed in the context of general first-order systems with interaction kernels in $W^{-1,\infty}$, a framework that includes the point vortex model approximating the two-dimensional Navier–Stokes equation. We also refer to [42], where the mean-field limit and quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels are established via the modulated energy method. In contrast to the relative entropy approach, which operates at the level of the joint law of the particle system, the modulated energy method is formulated directly in terms of the empirical measure associated with the particles.
Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results. In [2] and [3], the strategies given in [23] and [42], were adapted to get quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes. Additionally, [20], by means of a logarithmic Sobolev inequality, derived uniform in time propagation of chaos for the 2D vortex model, adapting the strategy in [23]. Alternatively, with sharp estimates, global in time derivation of that model, based on relaxation estimates for the limiting equation and inequalities of Riesz modulated energy, was derived by [12], see also [41],[11] and [31], for the modulated energy method. Regarding sub-Coulombian kernels, [6] obtained the derivation of the mean-field approximation for Landau-like equations, while [40] obtained global-in-time mean-field convergence for gradient (conservative) diffusive flows. By extending the BBGKY hierarchy given in [30], for singular kernels, [47] obtained the optimal convergence rate for the relative entropy between the k-particle marginals and the corresponding tensorized law for weakly interacting diffusions with bounded interaction kernels.
With this in view we observe that, the relative entropy method has proven effective for deriving quantitative estimates, for singular kernels, under periodic boundary conditions, while the modulated energy method, which can be applied in the whole space, has primarily been developed for deterministic interaction. However, more recently, quantitative propagation of chaos for the 2D vortex model and the 2D Log Gas on the whole space was obtained in [48] and [49], [7], respectively. In these works, the authors exploit Li–Yau-type estimates and Hamilton-type heat kernel estimates to derive suitable bounds for the limiting solution and its derivatives up to second order, which allows them to extend the strategy given in [23], to $\mathbb{R}^{d}$.
Quantitative estimates in the whole space were derived, in the context of moderately interacting particles  [^2] The moderate interact model was introduced by Oelschlager in [34],[33] and [35]. with singular kernels - including the Biot-Savart, the Keller-Segel and the Dirac delta measure kernels - by [36] and [37], based on a semigroup approach introduced in [17], with further advancements in [8], [21], [38], and [27]. It is important to note that these works do not rely on the relative entropy method, but instead establish quantitative estimates in strong functional topologies. The connection between the time evolution of the relative entropy, of the joint distribution, and moderate interacting regime, on the whole space was carried out by [10], based on a combination between the relative entropy and the regularised $L^{2}$-estimate in [35], deriving a propagation of chaos result for the viscous porous medium equation.
But was in [39], on the torus, that the equation (1), was derived, for singular kernels and common noise, in the context of moderate interaction regime, by applying the relative entropy directly to the mollified empirical measure. Earlier, also the same domain, the results established in [36] were extended to the setting with environmental noise in [26]. It is worth highlighting, that quantitative entropy estimates for systems with individual and common noise, as shown in [43] for incompressible the Navier-Stokes equations, [9] for the Hegselmann-Krause model, and [32] for mean-field systems with bounded kernels. For more results, see [18], [15], [16], and [28].
In contrast to the aforementioned works, our application of the relative entropy method provides pathwise estimates, deriving bounds directly at the trajectory level of the particle system. In those works, the estimates are obtained at the level of the joint law.
In order to present our results, this work is organized as follows. In the next two subsections, 1.1 and 1.2, respectively, we introduce the notations and assumptions used throughout the text, setting the stage for the subsequent discussions. In the subsection 1.3, we present our main results. The section 2 is dedicated to the proof of our main results. To ease the presentation, the Appendix, provides the corresponding suitable notion of a solution for the limiting equation, as well as some inequalities used throughout the proofs.
1.1 Notations
For $d\geq 1$, the space of probability density functions on $\mathbb{R}^{d}$ is denoted by $\mathcal{P}(\mathbb{R}^{d})$. For a normed vector space $U$, we denote its norm by $\|\cdot\|_{U}$, except for $L^{p}$ spaces, where we write $\|\cdot\|_{p}$, for $p\geq 1$. The Kantorovich-Rubinstein metric reads, for any two probability measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$,
$\|\mu-\nu\|_{0}\doteq\sup\left\{\int_{\mathbb{R}^{d}}\phi\,d(\mu-\nu)\,;\|\phi\|_{\infty},\|\phi\|_{\text{Lip}}\leq 1\right\}.$ (6)
Let $f$ and $g$ be positive probability density functions on $\mathbb{R}^{d}$. The relative entropy (or Kullback–Leibler divergence) of $f$ with respect to $g$ is defined as
$\mathcal{H}\left(f|g\right)\doteq\int_{\mathbb{R}^{d}}f(x)\ln{\left(\frac{f(x)}{g(x)}\right)}\,dx.$
Also the Fisher information of $f$ with respect to $g$ is given by
$\mathcal{I}\left(f|g\right)\doteq\int_{\mathbb{R}^{d}}f(x)\left|\nabla\ln{\frac{f(x)}{g(x)}}\right|^{2}\,dx=\int_{\mathbb{R}^{d}}\frac{(g(x))^{2}}{f(x)}\left|\nabla\left(\frac{f(x)}{g(x)}\right)\right|^{2}\,dx.$
1.2 Assumptions
  1. $(\mathbf{A}^{V})$
    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)=N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V:\mathbb{R}^{d}\to\mathbb{R}$ satisfying
    $\left|\nabla V\right|\leq C_{d}V,$ (7)
    and for $\alpha>d/2$
    $V(y)\leq C_{d}(1+|y|^{2})^{-\alpha}.$ (8)
  2. $(\mathbf{A}^{\nabla\cdot K})$
    It holds $\nabla\cdot K=0$ and there exists $K_{0}\in L^{\infty}$, with $\|K_{0}\|_{\infty}\leq 1/4$, such that
    $K=\nabla\cdot K_{0}.\,$ (9)
  3. $(\mathbf{A}^{\sigma})$
    The coefficient $\sigma:[0,T]\to\mathbb{R}^{d\times d}$ is measurable and bounded.
  4. $(\mathbf{A}^{\rho_{0}})$
    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,\infty}\left(\mathbb{R}^{d}\right)$ and
    $\left|\nabla\ln{\rho_{0}}(x)\right|^{2}\leq\widetilde{C}_{1}(1+|x|^{2}),$ (10)
    $\left|\nabla^{2}\ln{\rho_{0}}(x)\right|\leq\widetilde{C}_{2}(1+|x|^{2})$ (11)
    and
    $\rho_{0}(x)\leq\widetilde{C}_{3}\exp{\left(-\widetilde{C}_{3}^{-1}|x|^{2}\right)}.$ (12)
  5. $(\mathbf{A}^{X_{0}})$
    The initial particle system $(X_{0}^{i,N})_{i=1}^{N}$, $N\in\mathbb{N}$ is a sequence of i.i.d with law $\rho_{0}$.
Remark 1.
For an example of mollifier $V:\mathbb{R}^{d}\to\mathbb{R}$ satisfying $\left(\mathbf{A}^{V}\right)$, take
$V(y)\doteq\overline{C}_{d}\exp{\left(-\sqrt{1+|y|^{2}}\right)}$
with $\overline{C}_{d}\doteq\left(\int_{\mathbb{R}^{d}}\exp{\left(-\sqrt{1+|y|^{2}}\right)}\,dy\right)^{-1}$.
Remark 2.
The Biot-Savart kernel in dimension two is given by
$K(x)=\frac{1}{2\pi}\frac{x^{\perp}}{|x|^{2}}$ (13)
and then, it is divergence in the sense of distributions, of a $L^{\infty}$ matrix $K_{0}$, namely $K=\nabla\cdot K_{0}$, with
$K_{0}(x_{1},x_{2})=\left(\frac{1}{2\pi}\arctan{\left(\frac{x_{1}}{x_{2}}\right)}\right)I.$ (14)
It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$.
1.3 Statement of the main results
The first result, whose proof is postponed to the Appendix, establishes that the limiting equation is well posed in the appropriate functional framework. Indeed, it provides the regularity properties required in the proof of the main results.
Theorem 1.
Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. Then, there exist $T_{1}>0$, $C_{3}>0$, such that if $\widetilde{C}_{3}<C_{3}$,
  • there exists a pathwise solution of (1) with initial condition $\rho_{0}$, such that $\rho\in L^{\infty}\left([0,T_{1}];W^{2,1}\cap W^{2,\infty}(\mathbb{R}^{d})\right)$, $\mathbb{P}$-a.s.;
  • the solution $\rho$ verifies, $\mathbb{P}$-a.s., that
    $\left|\nabla\ln{\rho}(t,x)\right|^{2}\leq C_{1}(1+|x-X_{t}|^{2}),$ (15)
    $\left|\nabla^{2}\ln{\rho}(t,x)\right|\leq C_{2}(1+|x-X_{t}|^{2})$ (16)
    and
    $\rho(t,x)\leq\exp{(-2\pi|x-X_{t}|^{2})},$ (17)
    where
    $dX_{t}=\sigma_{t}dB_{t},$ (18)
    for all $t\in[0,T_{1}]$ and constants $C_{1},C_{2}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
We are now able to extend what was known before for bounded kernels in the whole space as in [32] and in [39], for sub-Coulombian kernels, under periodic boundary conditions. Our main result can be stated as follows.
Theorem 2.
Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$ and let $\rho$ be a solution of (1), given by Theorem 1. In addition, let the dynamics of the particle system be given by (2) and
$\lim_{N\to\infty}N^{\theta}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.$
where
$\theta=\min\left(1-\beta(2+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
with $\delta>0$, such that $\theta>0$, $d\geq 1$ and $\beta\in\left(0,\frac{1}{2\big[1+\frac{1}{d}\big]}\right)\cap\left(0,\frac{1}{\big[2+\frac{2}{d}+2\alpha\big]}\right)$.
Then, there exists a time $T\in(0,T_{1})$, such that
$\lim_{N\to\infty}N^{\theta}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})=0,\,\,\mathbb{P}-a.s.$
As a corollary of the previous theorem, we establish propagation of chaos for the marginals of the empirical measure of the particle system. In the sense of [1, Section 8.3], it corresponds to convergence of the empirical measure with respect to the Kantorovich–Rubinstein metric, given in (6).
Corollary 1.
Under assumptions of Theorem 2, we have
$\lim_{N\to\infty}N^{\theta-\delta}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}=0,\,\,\,\mathbb{P}-a.s.$
Remark 3.
We observe that, compared with previous works on moderate particle systems without common noise, we lose the order of convergence. This is because the estimation of the martingale term in our work is of order $\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)$, while in [36] is the order $\frac{1}{2}-\frac{\beta}{2}$, see Theorem 1.3 in [36]. When environmental noise is taken into account, the overall framework remains comparable to that of [26]. For the same reason, the range of the parameter $\beta$ is smaller in our work. However, in the present work we obtain an improved convergence rate compared to that established in [39].
2 Proofs of main results
In this section we present the proofs of our main results.
In Subsection 2.1, by applying the Itô’s formula, such as in [39], we derive an evolution equation for the relative entropy functional of the regularized empirical measure with respect to the solution of (1).
In Subsection 2.2, we deal with the quadratic variation terms arising from the application of Itô’s formula.
In Subsection 2.3, we address the nonlinear contributions appearing in the evolution equation for the relative entropy functional obtained in the first step.
Finally, in Subsections 2.4 and 2.5, combining the previous estimates with Gronwall’s lemma and suitable localization techniques, we close the argument and complete the proof.
2.1 Time evolution of the relative entropy
By applying Itô’s formula, following the same arguments developed in Section 2 of [39], we derive the corresponding evolution equation for the entropy functional between the regularized empirical measure $\rho^{N}$ and the solution $\rho$ of (1):
$\mathcal{H}(\rho_{t}^{N}|\rho_{t})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $=I_{t}+II_{t}+III_{t}+M_{t}^{N}+IV_{t},$ (19)
where
$I_{t}$ $\overset{}{=}\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho_{s}^{N}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)[K\ast\rho_{s}-K\ast\rho_{s}^{N}]\,dx\,ds$
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds,$ (20)
$II_{t}+IV_{t}$ $=\frac{1}{2N^{2}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{1}{\rho_{s}^{N}}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx,$ (21)
$III_{t}$ $=-\frac{1}{2}\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds,$ (22)
and
$M_{t}^{N}$ $=\int_{\mathbb{R}^{d}}\int_{0}^{t}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dB_{s}\,dx$
$+\int_{\mathbb{R}^{d}}\int_{0}^{t}\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dB_{s}\,dx$
$+\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\nabla V^{N}(x-X_{s}^{i})\,dW^{i}_{s}\,dx$
$-\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\nabla V^{N}(x-X_{s}^{i})\,dW^{i}_{s}\,dx$
$+\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{\rho_{s}^{N}}{\rho_{s}}\sigma^{\top}_{s}\nabla\rho_{s}\,dB_{s}\,dx.$ (23)
2.2 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
We now derive estimates for the quadratic variation terms appearing in (19). The main novelty consists in introducing a suitable stopping time in order to handle the integral over the whole space arising in this term. It is important to emphasize that the argument used in [39], which relies on periodic boundary conditions, is no longer available in the present setting.
First, taking into account Assumption $(\mathbf{A}^{V})$ in (21), we have
$II_{t}+IV_{t}$ $\leq\frac{N^{2\beta/d}}{2N^{2}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{1}{\rho_{s}^{N}}|V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx.$ (24)
We introduce the following stopping time:
$\tau^{N}=\inf\left\{t\geq 0;\exists i\in\{1,...,N\},|X_{t}^{i,N}|\geq N^{\beta}\right\}\wedge T.$ (25)
By decay of $\rho_{0}$ in (12), along with Borel-Cantelli Lemma, almost surely, there exists $N_{0}=N_{0}(\omega)$ such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$.
Indeed, setting
$\Lambda^{N}\doteq\left\{\exists i\in\{1,...,N\},|X_{0}^{i,N}|\geq N^{\beta}\right\}=\bigcup_{i=1}^{N}\left\{|X_{0}^{i,N}|\geq N^{\beta}\right\}$ (26)
by Assumption $(\mathbf{A}^{X_{0}})$, we derive
[…]
$=\sum_{i=1}^{N}\int_{|x|\geq N^{\beta}}\rho_{0}(x)\,dx$
$\leq C_{d}\sum_{i=1}^{N}(N^{\beta})^{d-2}e^{-\widetilde{C}_{3}^{-1}(N^{\beta})^{2}}$
$\leq C_{d}N(N^{\beta})^{d-2}e^{-\widetilde{C}_{3}^{-1}(N^{\beta})^{2}}$ (27)
where in the second inequality we are using (12) to get
$\int_{|x|\geq N^{\beta}}\rho_{0}(x)dx$ $\leq\widetilde{C}_{3}\int_{|x|\geq N^{\beta}}e^{-\widetilde{C}_{3}^{-1}|x|^{2}}dx$
$=C_{d}\int_{r=N^{\beta}}^{\infty}e^{-\widetilde{C}_{3}^{-1}r^{2}}r^{d-1}dr\leq C_{d}(N^{\beta})^{d-2}e^{-\widetilde{C}_{3}^{-1}(N^{\beta})^{2}}.$ (28)
It follows by Borel-Cantelli Lemma,
$\mathbb{P}\left(\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}\Lambda^{N}\right)=0.$ (29)
With this in view for all $\omega\in\Lambda_{0}$, with
$\Lambda_{0}\doteq\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}$ (30)
there exists $N_{0}=N_{0}(\omega)$, such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$, since $t\mapsto X_{t}^{i,N}(\omega)$ is continuous, almost surely.
Let $s\in(0,\tau^{N})$. We observe that,
$I^{N}\doteq\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|^{2}}{\rho^{N}_{s}(x)}\,dx$ $=\int_{|x|\leq 2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|^{2}}{\rho^{N}_{s}(x)}\,dx$
$+\int_{|x|>2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|^{2}}{\rho^{N}_{s}(x)}\,dx.$ (31)
Now, if $|x|\leq 2|X_{s}^{i,N}|$, then $|x|\leq 2N^{\beta}$ for $s\in(0,\tau^{N})$, and since $V^{N}\leq N^{\beta}$, we derive for $\alpha>d/2$
$V^{N}(x-X_{s}^{i,N})\leq N^{\beta}\left(1+|x|^{2}\right)^{\alpha-\alpha}\leq C_{\alpha}N^{2\alpha\beta+\beta}\left(1+|x|^{2}\right)^{-\alpha}.$ (32)
On the other hand, if $|x|>2|X_{s}^{i,N}|$ we have $|x-X_{s}^{i,N}|>\frac{|x|}{2}$ and by decay of $V$, given in (8), we get
[…]
$\leq C_{\alpha}N^{\beta}\left(1+\frac{|N^{\frac{\beta}{d}}x|^{2}}{4}\right)^{-\alpha}\leq C_{\alpha}N^{\beta}\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}.$ (33)
Therefore by (31), (32) and (33), we find
$I^{N}$ $\leq C_{\alpha}\int_{|x|\leq 2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\left(N^{2\alpha\beta+\beta}\left(1+|x|^{2}\right)^{-\alpha}\right)\,dx$
$+C_{\alpha}\int_{|x|>2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\left(N^{\beta}\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\right)\,dx$
$\leq C_{\alpha}N^{2\alpha\beta+\beta}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\left(1+|x|^{2}\right)^{-\alpha}\,dx$
$+C_{\alpha}N^{2\alpha\beta+\beta}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\,dx$
$=C_{\alpha}N^{2\alpha\beta+\beta}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx,$ (34)
with
$\phi(x)\doteq\left(\left(1+|x|^{2}\right)^{-\alpha}+\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\right)$ (35)
Hence by (34) in (24) we obtain
$II_{t\wedge\tau^{N}}+IV_{t\wedge\tau^{N}}$ $=\frac{1}{2N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{1}{\rho_{s}^{N}(x)}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,dx\,ds$
$\overset{}{\leq}\frac{N^{\beta}N^{\frac{2\beta}{d}}}{2N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{1}{\rho_{s}^{N}(x)}|V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
$\overset{\alpha>d/2}{\leq}C_{\alpha,T}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}.$ (36)
2.3 Estimates for $I_{t}$: the nonlinear terms
We now address the nonlinear terms in (19). The main novelty is to use Donsker-Varadhan inequality (79), in the context of moderate interactions, along with the dissipation that comes from of Fisher information and Assumption $(\mathbf{A}^{\nabla\cdot K})$, to tackle with the singularity of kernel $K$.
First, we set
$I_{t}$ $\overset{}{=}\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho_{s}^{N}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)[K\ast\rho_{s}-K\ast\rho_{s}^{N}]\,dx\,ds$
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
$\doteq I^{1}_{t}-I^{2}_{t},$ (37)
where
$I^{1}_{t}$ $=-\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\rho_{s}^{N}\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\nabla\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
$\doteq-I^{1,1}_{t}-I^{1,2}_{t}.$ (38)
Now by $\epsilon$-Young inequality and convolution inequality, we deduce
$I_{t}^{1,2}$ $=\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\nabla\rho_{s}\left(\frac{\sqrt{\rho_{s}^{N}}}{\rho_{s}}\right)\left(\frac{\rho_{s}}{\sqrt{\rho_{s}^{N}}}\right)\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\left|K_{0}\ast(\rho_{s}-\rho_{s}^{N})\right|^{2}\,dx\,ds$
$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\|K_{0}\|\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\,dx\,ds.$ (39)
We observe that by Donsker-Varadhan inequality (79), for $\eta>>1$
[…]
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$
which implies by Csiszár-Kullback-Pinsker inequality (78) and (39)
$I_{t}^{1,2}$ $\leq\epsilon\int_{0}^{t}\mathcal{I}(\rho^{N}_{s}|\rho_{s})\,ds+C_{\epsilon,\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds.$ (40)
Regarding term $I_{t}^{1,1}$ in (38), by convolution inequality, we find
$I_{t}^{1,1}$ $=\int_{0}^{t}\|K_{0}\|_{\infty}\|\rho^{N}_{s}-\rho_{s}\|_{1}\int_{\mathbb{R}^{d}}\left|\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right|\rho^{N}_{s}\,dx\,ds.$ (41)
In addition, if $C_{1,2}\doteq\max{(C_{1},C_{2})}$ since
[…]
by Donsker-Varadhan inequality (79), we obtain
[…]
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{\eta>>1}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\eta\ln{(1)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s}),$ (43)
for $\eta>>1$.
Therefore, by (43) in (41), we get
$I_{t}^{1,1}$ $\overset{}{\leq}C_{\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds.$ (44)
[…]
$I_{t}^{1}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\epsilon\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds.$ (45)
Now, we will focus on the term $I_{t}^{2}$ in (37). First, we recall that by Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$, $K=\nabla\cdot K_{0}$ with $K_{0}\in L^{\infty}$. Hence, we arrive at
$\left|K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right|$ $=\left|K_{0}\ast\nabla\rho_{s}^{N}(\cdot)-K_{0}\ast\nabla\rho_{s}^{N}(x)\right|$
$\leq\int_{\mathbb{R}^{d}}\left|K_{0}(\cdot-y)-K_{0}(x-y)\right||\nabla\rho_{s}^{N}(y)|\,dy$
$\leq\int_{\mathbb{R}^{d}}2\|K_{0}\|_{\infty}|\nabla\rho_{s}^{N}(y)|\,dy$
$=2\|K_{0}\|_{\infty}\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}(y)|\,dy.$ (46)
Additionally, $\rho_{s}^{N}\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, $\mathbb{P}$-a.s. and then by Holder’s inequality we obtain
$\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\,dx$ $=\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\frac{\sqrt{\rho_{s}^{N}}}{\sqrt{\rho_{s}^{N}}}\,dx$
$\overset{\text{Holder}}{\leq}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}^{d}}\rho_{s}^{N}\,dx\right)^{\frac{1}{2}}$
$\overset{}{=}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}.$ (47)
Now from Leibniz’s rule we deduce, for $\eta>>1$
$\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx$ $=\int_{\mathbb{R}^{d}}\frac{\left|\nabla\left(\rho_{s}\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
$\leq 2\int_{\mathbb{R}^{d}}\frac{\left|\nabla\rho_{s}\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx+2\int_{\mathbb{R}^{d}}\frac{\left|\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
$\overset{}{\leq}2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\frac{\left|\nabla\rho_{s}\right|^{2}}{\rho_{s}^{2}}\,dx+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx$
$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (48)
where in the last inequality, we are using Donsker-Varadhan inequality (79):
[…]
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{\eta>>1}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\eta\ln{(1)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s}).$
[…]
$\left|\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\right|$ $\leq 2\|K_{0}\|_{\infty}\left\langle S_{s}^{N},V^{N}(x-\cdot)\right\rangle$
$\times\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}$
$\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}},$
which implies, by $\epsilon$-Young and Holder inequalities,
$I^{2}_{t}$ $\doteq\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\,dx\,ds$
$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})\right)^{\frac{1}{2}}\,dx\,ds$
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\,dx\,ds$
$\overset{\text{Young}}{\leq}\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds$
$+\sqrt{2}/2\int_{0}^{t}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\,dx\,ds$
$\overset{\text{Holder}}{\leq}(\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds.$ (49)
[…]
$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds.$ (50)
2.4 Application of Gronwall’s Lemma
[…]
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds$
$\overset{}{+}(2\epsilon+\sqrt{2}/2)\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
$-\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
$+C_{\alpha,T}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}$
$+M_{t}^{N}.$ (51)
With the same computations as in [39], for all $\delta\in(0,\theta_{1})$ there exists a random variable $A_{0}$, with finite moments such that
$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{1}+\delta},$
which allows us to express (51), for $\epsilon<<1$, as
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+C_{\alpha,T}N^{-\theta_{2}}+A_{0}N^{-\theta_{1}+\delta},$ (52)
with $\theta_{2}\doteq 1-\beta(2+\frac{2}{d}+2\alpha)$.
From Gronwall’s Lemma in (52), recalling (26), (30) and (29), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}\right)$ (53)
with
$\theta\doteq\min\left(1-\beta(2+\frac{2}{d}+2\alpha);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta\right).$
2.5 End of the proof: Removing the stopping time
We now remove the stopping time $\tau^{N}$ introduced in (53) by exploiting the probabilistic framework provided by Assumptions $\left(\mathbf{A}^{\rho_{0}}\right)$ and $\left(\mathbf{A}^{X_{0}}\right)$. These assumptions ensure sufficient control on the initial data and on the particle system, allowing us to pass from the localized estimates to global-in-time bounds with high probability, and finally eliminate the stopping time.
First, since $K=\nabla\cdot K_{0}$, $K_{0}\in L^{\infty}$ and $V^{N}=N^{\beta}V(N^{\beta/d}\cdot)$, by (2), we have
$\sup_{t\in[0,T]}|X_{t}^{i,N}|$ $\leq|X_{0}^{i,N}|+T\|\nabla V\|_{1}\|K_{0}\|_{\infty}N^{\beta}$
$+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sqrt{2}dW_{s}^{i,N}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|,$ (54)
which implies for $\widetilde{C}_{K_{0},V}\doteq(1-T\|\nabla V\|_{1}\|K_{0}\|_{\infty})$
$\{\tau^{N}<T\}$ $\subset\bigcup_{i=1}^{N}\left\{\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
$=\left\{\exists i;\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
$\subset\left\{\exists i;|X_{0}^{i,N}|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sqrt{2}dW_{s}^{i,N}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|\geq\widetilde{C}_{K_{0},V}N^{\beta}\right\}$
$\subset L_{1}^{N}\cup L_{2}^{N}\cup L_{3}^{N},$ (55)
with
$L_{1}^{N}\doteq\left\{\exists i;|X_{0}^{i,N}|\geq C_{K_{0},V}N^{\beta}\right\},$ (56)
$L_{2}^{N}\doteq\left\{\exists i;\sup_{t\in[0,T]}\left|\int_{0}^{t}dW_{s}^{i,N}\right|\geq C_{K_{0},V}N^{\beta}\right\}$ (57)
and
$L_{3}^{N}\doteq\left\{\exists i;\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|\geq C_{K_{0},V}N^{\beta}\right\}$ (58)
with $C_{K_{0},V}\doteq(1-T\|\nabla V\|_{1}\|K_{0}\|_{\infty})/3\sqrt{2}>0$, for $T<(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})$.
We note that, with the same computations like those led (27), we deduce
$\mathbb{P}\left(L_{1}^{N}\right)$ $\overset{}{\leq}N(C_{K_{0},V}N^{\beta})^{d-2}e^{-(C_{K_{0},V}N^{\beta})^{2}}$ (59)
We now take care about the $L_{2}^{N}$. Indeed, we note that $W_{t}^{i,N}=\int_{0}^{t}dW_{s}^{i,N}$, and since
$|W_{t}^{i,N}|^{2}=\sum_{j=1}^{d}(W_{t}^{i,N,j})^{2},$
if $|W_{t}^{i,N}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}.$ Therefore,
$\left\{\sup_{t\leq T}|W_{t}^{i,N}|\geq a\right\}\subset\bigcup_{j=1}^{d}\left\{\sup_{t\leq T}|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}\right\},$ (60)
we obtain
[…]
$\leq\sqrt{\frac{T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}$
$\times\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2T)\right)},$ (61)
where in the second inequality, we are using (8.3)’ in Chapter 2 of [25].
Now we will derive an analogous estimate for $L_{3}^{N}$. First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$. Since $\sigma$ is bounded, the function defined by $t\mapsto\int_{0}^{t}\sigma^{2}_{s}\,ds$ is continuous and increasing. Hence, recalling (18), by Theorem 4.6, in Chapter 3 of [25], for all $j\in\{1,...,d\}$, we get
$\sup_{t\leq T}|X_{t}^{j}|=\sup_{t\leq T}|\widetilde{B}^{j}_{\langle X\rangle_{t}}|\leq\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j}_{t}|,$ (62)
where $(\widetilde{B}_{t})_{t\geq 0}$ is the time-change, which in particular, is a one dimensional standard Brownian motion. In addition, since
$|X_{t}|^{2}=\sum_{j=1}^{d}|X_{t}^{j}|^{2}$
for
$X_{t}^{j}=\int_{0}^{t}\sigma_{s}\,dB_{s}^{j},$
if $|X_{t}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|X_{t}^{j}|\geq\frac{a}{\sqrt{d}}.$ Therefore, for all $a>0$, we find
[…]
It follows that, by (62) and (63)
[…]
$\leq\sqrt{\frac{\|\sigma\|_{\infty}^{2}T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}$
$\times\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2\|\sigma\|_{\infty}^{2}T)\right)},$ (64)
where in the second inequality, we are using (8.3)’ in Chapter 2 of [25].
Hence, by (59), (61) and (64) in (58), we find
$\sum_{N\in\mathbb{N}}\mathbb{P}\left(\tau^{N}<T\right)<\infty$ (65)
and by Borel-Cantelli Lemma, we arrive at
$\mathbb{P}\left(\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}\{\tau^{N}<T\}\right)=0.$ (66)
Finally, by (30) and (66), for $T<\min{\left(T_{1};(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})\right)}$ for all $\omega\in\Lambda$, with
$\Lambda\doteq\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}\right)\cap\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}\{\tau^{N}=T\}\right),$ (67)
there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)=T$ and (53) holds. It follows that, almost surely
$\lim_{N\to\infty}N^{\theta}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})=0,$
which conclude the proof of Theorem.
3 Appendix
This appendix contains the proof of Theorem 1 together with several auxiliary inequalities used in the previous sections. We begin by establishing Theorem 1, and subsequently derive the additional estimates required throughout the paper.
3.1 Proof of Theorem 1
3.1.1 Solving (1) for $\sigma=0$
When $\sigma=0$, a solution for (1) in the sense of Theorem 1, with exception of (17), was constructed in Sections 3 and 4 of [48] and [49]. With this in view, we only need to verify that, it is possible to take $\tilde{C}_{3}$ sufficiently small, in order to get that Gaussian decay.
Proof of (17).
By Theorem 3 in [13] along with (12), we obtain there exists $C>0$ such that for all $(t,x)\in[0,T_{1}]\times\mathbb{R}^{d}$
$\bar{\rho}(t,x)\leq C\widetilde{C}_{3}\int_{\mathbb{R}^{d}}\frac{1}{t^{d/2}}\exp{\left(-\frac{\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}}{8t\widetilde{C}_{3}}\right)}dy.$ (68)
Now since
$|x-y|^{2}=|x|^{2}-2x\cdot y+|y|^{2}.$
we find
$\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}$ $=\widetilde{C}_{3}|x|^{2}-2\widetilde{C}_{3}x\cdot y+\widetilde{C}_{3}|y|^{2}+8t|y|^{2}$
$=\widetilde{C}_{3}|x|^{2}-2\widetilde{C}_{3}x\cdot y+(\widetilde{C}_{3}+8t)|y|^{2}.$ (69)
Additionally, we get
$(\widetilde{C}_{3}+8t)|y|^{2}-2\widetilde{C}_{3}x\cdot y.=(\widetilde{C}_{3}+8t)\left(|y|^{2}-2\frac{\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}x\cdot y\right).$
and since
$|y|^{2}-2ax\cdot y=|y-ax|^{2}-a^{2}|x|^{2},$
with
$a=\frac{\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}.$
we arrive at
$\Big((\widetilde{C}_{3}+8t)|y|^{2}-2\widetilde{C}_{3}x\cdot y\Big)=(\widetilde{C}_{3}+8t)|y-ax|^{2}-\frac{\widetilde{C}_{3}^{2}}{\widetilde{C}_{3}+8t}|x|^{2}.$ (70)
Therefore by (69) and (70) we derive
$\exp\left(-\frac{\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}}{8t\widetilde{C}_{3}}\right)=\exp{\left(-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}\right)}\exp\left(-\frac{\widetilde{C}_{3}+8t}{8t\widetilde{C}_{3}}|y-ax|^{2}\right).$ (71)
and letting z=yaxz=y-ax we have
$\bar{\rho}(x,t)$ $\leq C\widetilde{C}_{3}\frac{1}{t^{d/2}}e^{-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}}\int_{\mathbb{R}^{d}}\exp{\left(-\frac{\widetilde{C}_{3}+8t}{8t\widetilde{C}_{3}}|z|^{2}\right)}dz$
$\leq C\widetilde{C}_{3}\frac{1}{t^{d/2}}e^{-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}}\left(\frac{\pi 8t\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}\right)^{d/2}$
$\leq C\widetilde{C}_{3}(\pi 8\widetilde{C}_{3})^{d/2}\frac{1}{(\widetilde{C}_{3}+8t)^{d/2}}\exp{\left(-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}\right)}$ (72)
since
$\int_{\mathbb{R}^{d}}e^{-a|z|^{2}}dz=\left(\frac{\pi}{a}\right)^{d/2}.$
It follows that taking
$\widetilde{C}_{3}<C_{3}\doteq\min\left(\frac{1}{2\pi};\frac{1}{(8\pi)^{d/2}C}\right),$ (73)
for $t\in(0,T]$ such that
$t<T_{1}\doteq\min\left(\frac{\widetilde{C}_{3}}{8};\frac{8(1-2\pi\widetilde{C}_{3})}{2\pi}\right)$ (74)
we end up with
$\bar{\rho}(x,t)\leq\exp{(-2\pi|x|^{2})}.$ (75)
3.1.2 Solving (1) for $\sigma\neq 0$
We assume $\tilde{\rho}$ solves the deterministic PDE
$\partial_{t}\tilde{\rho}-\Delta\tilde{\rho}+\nabla\cdot(\tilde{\rho}_{t}K\ast\tilde{\rho}_{t})=0$ (76)
Now we define $\rho(t,x)\doteq\tilde{\rho}(t,x-X_{t})$, where
$dX_{t}=\sigma_{t}dB_{t},$ (77)
Hence, by applying the Itô’s formula, since that $\sigma$ does not depends on spatial variable, we find that $\rho$ solves (1), in the sense of Theorem 1.
3.2 Inequalities
In this section, we present several inequalities used throughout the text. The first one is the classical Csiszár-Kullback-Pinsker inequality, which is a fundamental result in information theory. It quantifies how control of the relative entropy implies control of the total variation distance.
Lemma 3 (Csiszár-Kullback-Pinsker inequality).
It holds for $f,g\in\mathcal{P}\left(\mathbb{R}^{d}\right)$,
$\|f-g\|_{1}^{2}\lesssim\mathcal{H}(f|g).$ (78)
Now, we recall the Donsker–Varadhan inequality (see [23]) which follows from the variational representation of the relative entropy. It provides a tool to estimate the expectation of a test function under a non-factorized probability density and, in particular, allows for a change of measure between two probability measures.
Lemma 4 (Donsker-Varadhan Inequality).
Let $f,g\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, $\eta>0$ and a function $\Phi$ on $\mathbb{R}^{d}$ such that $g\exp{\left(\Phi/\eta\right)}\in L^{1}$. Then, we have
$\int_{\mathbb{R}^{d}}f\Phi\,dx\leq\eta\mathcal{H}\left(f|g\right)+\eta\ln{\left(\int_{\mathbb{R}^{d}}g\exp{\left(\Phi/\eta\right)}\,dx\right)}.$ (79)
Acknowledgements
A. B. de Souza is partially supported by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code $001$.

### v2

@@
-Quantitative propagation of chaos for 2D stochastic vortex model on the whole space under moderate interactions
+arXiv:2602.22038v2[math.AP] 27 Feb 2026
 Abstract
-We derive the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in the sense of relative entropy.
-The main novelty is the application of the Donsker–Varadhan inequality in the context of moderately interacting particles to handle the nonlinearity, as well as the use of localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms.
+We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in the terms of relative entropy.
+The main novelties lie in the application of the Donsker–Varadhan inequality within the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with a probabilistic data setting, in order to derive estimates for the quadratic variation terms.
 Moreover, we prove the existence of a suitable solution to the aforementioned model.
@@
 where $\theta$ is an explicit positive parameter, by (4) and (5), we have, in particular, propagation of chaos for the marginals of the empirical measure of the particle system, see [1, Section 8.3] and [45].
-In order to derive (5), we apply for the first time in the context of moderately interacting particles, the Donsker–Varadhan inequality together with the dissipation provided by the Fisher information, to handle the nonlinearity.
+In the present work, we apply for the first time in the context of moderately interacting particles, the Donsker–Varadhan inequality together with the dissipation provided by the Fisher information, to handle the nonlinearity.
 In addition, we employ localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms in the whole space, as the estimate established in [39], under periodic boundary conditions, does not apply in the present framework.
@@
 It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$.
+Remark 3.
+Assumption $(\mathbf{A}^{\rho_{0}})$ was formulated in [7, 48, 49], where, relying on Li–Yau–type gradient bounds and Hamilton–type heat kernel estimates, the authors use it to control the solution and its derivatives up to second order.
+It allows the relative entropy approach of [23] to be extended from the periodic framework to the whole space.
 1.3 Statement of the main results
@@
     and
-    $\rho(t,x)\leq\exp{(-2\pi|x-X_{t}|^{2})},$ (17)
+    $\rho(t,x)\leq 2\exp{(-2\pi|x-X_{t}|^{2})},$ (17)
     where
@@
 In addition, let the dynamics of the particle system be given by (2) and
-$\lim_{N\to\infty}N^{\theta}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.$
+$\mathcal{H}(\rho_{0}^{N}|\rho_{0})\leq C,\,\,\,\mathbb{P}-a.s.$
 where
-$\theta=\min\left(1-\beta(2+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
-with $\delta>0$, such that $\theta>0$, $d\geq 1$ and $\beta\in\left(0,\frac{1}{2\big[1+\frac{1}{d}\big]}\right)\cap\left(0,\frac{1}{\big[2+\frac{2}{d}+2\alpha\big]}\right)$.
-Then, there exists a time $T\in(0,T_{1})$, such that
-$\lim_{N\to\infty}N^{\theta}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})=0,\,\,\mathbb{P}-a.s.$
+$\theta=\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
+with $\delta>0$, such that $\theta>0$, $d\geq 1$ and $\beta\in\left(0,\frac{1}{2\big[1+\frac{1}{d}\big]}\right)\cap\left(0,\frac{1}{\big[1+\frac{2}{d}+2\alpha\big]}\right)$.
+Then, there exist a time $T\in(0,T_{1})$ and a constant $C>0$, such that, almost surely for $N>>1$,
+$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (19)
 As a corollary of the previous theorem, we establish propagation of chaos for the marginals of the empirical measure of the particle system.
@@
 Under assumptions of Theorem 2, we have
-$\lim_{N\to\infty}N^{\theta-\delta}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}=0,\,\,\,\mathbb{P}-a.s.$
+$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (20)
 Remark 4.
@@
 Hence by (36) in (26) we obtain
-$II_{t\wedge\tau^{N}}+IV_{t\wedge\tau^{N}}$ $=\frac{1}{2N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{1}{\rho_{s}^{N}(x)}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,dx\,ds$
-$\overset{}{\leq}\frac{N^{\beta}N^{\frac{2\beta}{d}}}{2N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{1}{\rho_{s}^{N}(x)}|V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx$
-$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
-$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
-$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
-$\overset{\alpha>d/2}{\leq}C_{\alpha,T}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}.$ (36)
+[…]
+$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
+$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
+$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
+$\overset{\alpha>d/2}{\leq}C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}.$ (38)
 2.3 Estimates for $I_{t}$: the nonlinear terms
@@
 […]
-$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
-$\overset{\eta>>1}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\eta\ln{(1)}$
-$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s}),$ (43)
+$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
+$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$ (45)
 for $\eta>>1$.
 Therefore, by (45) in (43), we get
-$I_{t}^{1,1}$ $\overset{}{\leq}C_{\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds.$ (44)
+$I_{t}^{1,1}$ $\overset{}{\leq}C_{\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (46)
 […]
-$I_{t}^{1}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\epsilon\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds.$ (45)
+$I_{t}^{1}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\epsilon\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (47)
 Now, we will focus on the term $I_{t}^{2}$ in (39). First, we recall that by Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$, $K=\nabla\cdot K_{0}$ with $K_{0}\in L^{\infty}$. Hence, we arrive at
@@
 $\overset{}{\leq}2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\frac{\left|\nabla\rho_{s}\right|^{2}}{\rho_{s}^{2}}\,dx+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx$
-$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (48)
+$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (50)
 where in the last inequality, we are using Donsker-Varadhan inequality (82):
@@
 […]
-$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
-$\overset{\eta>>1}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\eta\ln{(1)}$
-$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s}).$
+$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
+$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}.$
 […]
@@
 $\times\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}$
-$\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}},$
+$\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}},$
 which implies, by $\epsilon$-Young and Holder inequalities,
 $I^{2}_{t}$ $\doteq\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
-$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\,dx\,ds$
+$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
 $\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})\right)^{\frac{1}{2}}\,dx\,ds$
 $+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\,dx\,ds$
+$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
 $\overset{\text{Young}}{\leq}\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds$
 $+\sqrt{2}/2\int_{0}^{t}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\,dx\,ds$
-$\overset{\text{Holder}}{\leq}(\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds.$ (49)
+$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\eta}$
+$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\eta}.$ (51)
 […]
-$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds.$ (50)
+$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (52)
 2.4 Application of Gronwall’s Lemma
+In this subsection, we conclude our estimates, by applying Gronwall’s inequality.
+Indeed, with the same computations as in [39], for all $\delta\in(0,\theta_{1})$ there exists a random variable $A_{0}$, finite almost surely, such that
+$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{1}+\delta},$
+with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
 […]
 $\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds$
-$\overset{}{+}(2\epsilon+\sqrt{2}/2)\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
+$\overset{}{+}(3\epsilon+\sqrt{2}/2)\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
 $-\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
-$+C_{\alpha,T}\frac{N^{\beta(2+\frac{2}{d}+2\alpha)}}{N}$
-$+M_{t}^{N}.$ (51)
-With the same computations as in [39], for all $\delta\in(0,\theta_{1})$ there exists a random variable $A_{0}$, with finite moments such that
-$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{1}+\delta},$
-which allows us to express (51), for $\epsilon<<1$, as
-$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+C_{\alpha,T}N^{-\theta_{2}}+A_{0}N^{-\theta_{1}+\delta},$ (52)
-with $\theta_{2}\doteq 1-\beta(2+\frac{2}{d}+2\alpha)$.
-From Gronwall’s Lemma in (52), recalling (26), (30) and (29), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
-$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}\right)$ (53)
+$+C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}$
+$+A_{0}N^{-\theta_{1}+\delta}$
+$+TC_{\eta}$ (53)
+and then for $\eta>>1$ and $\epsilon<<1$, we deduce
+$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{2}}+A_{0}N^{-\theta_{1}+\delta}+C,$ (54)
+with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$ and $\theta_{2}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$.
+Finally, from Gronwall’s Lemma in (54), recalling (28), (31) and (32), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
+$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (55)
 with
-$\theta\doteq\min\left(1-\beta(2+\frac{2}{d}+2\alpha);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta\right).$
+$\theta\doteq\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta\right).$
 2.5 End of the proof: Removing the stopping time
@@
 First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$.
-Since $\sigma$ is bounded, the function defined by $t\mapsto\int_{0}^{t}\sigma^{2}_{s}\,ds$ is continuous and increasing.
+Since $\sigma$ is bounded, for $j\in\{1,...,d\}$, the function defined by $t\mapsto\langle X^{j}\rangle_{t}$ is continuous and increasing.
 Hence, recalling (18), by Theorem 4.6, in Chapter 3 of [25], for all $j\in\{1,...,d\}$, we get
-$\sup_{t\leq T}|X_{t}^{j}|=\sup_{t\leq T}|\widetilde{B}^{j}_{\langle X\rangle_{t}}|\leq\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j}_{t}|,$ (62)
-where $(\widetilde{B}_{t})_{t\geq 0}$ is the time-change, which in particular, is a one dimensional standard Brownian motion. In addition, since
+$\sup_{t\leq T}|X_{t}^{j}|=\sup_{t\leq T}\left|\widetilde{B}^{j}_{\langle X^{j}\rangle_{t}}\right|\leq\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j}_{t}|,$ (64)
+where $(\widetilde{B}^{j}_{t})_{t\geq 0}$ is the time-change process, which in particular, is a one dimensional standard Brownian motion. In addition, since
 $|X_{t}|^{2}=\sum_{j=1}^{d}|X_{t}^{j}|^{2}$
-for
-$X_{t}^{j}=\int_{0}^{t}\sigma_{s}\,dB_{s}^{j},$
 if $|X_{t}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|X_{t}^{j}|\geq\frac{a}{\sqrt{d}}.$ Therefore, for all $a>0$, we find
@@
 $\sum_{N\in\mathbb{N}}\mathbb{P}\left(\tau^{N}<T\right)<\infty$ (67)
-and by Borel-Cantelli Lemma, we arrive at
+and then by Borel-Cantelli Lemma, we arrive at
 $\mathbb{P}\left(\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}\{\tau^{N}<T\}\right)=0.$ (68)
@@
 there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)=T$ and (55) holds. It follows that, almost surely
-$\lim_{N\to\infty}N^{\theta}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})=0,$
+$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (70)
 which conclude the proof of Theorem.
@@
 we end up with
-$\bar{\rho}(x,t)\leq\exp{(-2\pi|x|^{2})}.$ (75)
+$\bar{\rho}(x,t)\leq 2\exp{(-2\pi|x|^{2})}.$ (78)
 3.1.2 Solving (1) for $\sigma\neq 0$

### v3

@@
-arXiv:2602.22038v2[math.AP] 27 Feb 2026
+Quantitative entropy estimates for 2D stochastic vortex model on the whole space under moderate interactions
 Abstract
-We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in the terms of relative entropy.
-The main novelties lie in the application of the Donsker–Varadhan inequality within the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with a probabilistic data setting, in order to derive estimates for the quadratic variation terms.
-Moreover, we prove the existence of a suitable solution to the aforementioned model.
-MSC2010 subject classification: 49N90, 60H30, 60K35
+We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in terms of relative entropy.
+The main novelties lie in the application of the Donsker–Varadhan inequality in the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with the probabilistic data setting, in order to derive estimates for the quadratic variation terms.
+As an application of our entropy bounds, we derive a new energy estimate for the distance between the regularized empirical measure and the solution to the aforementioned model by combining the control of the Fisher information of the particle system, the Ladyzhenskaya inequality, and a nonlinear Grönwall lemma.
+Moreover, we prove the existence of a suitable solution to the limiting process.
+MSC2010 subject classification: 35R60, 35Q30, 82C22
 1 Introduction
-In this paper, we establish a quantitative derivation of the stochastic 2D vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
+In this work, we derive quantitative entropy estimates for the 2D stochastic vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
 $\mathrm{d}\rho_{t}=\Delta\rho_{t}\,dt+\frac{1}{2}D^{2}\rho_{t}(\sigma\sigma^{\top})_{t}\,dt-\nabla\cdot(\rho_{t}(K\ast\rho_{t}))\,dt-\nabla\rho_{t}\cdot\sigma_{t}\,dB_{t},$ (1)
@@
 $dX_{t}^{i,N}$ $=\frac{1}{N}\sum_{k=1}^{N}\left(K\ast V^{N}\right)\left(X_{t}^{i,N}-X_{t}^{k,N}\right)\,dt+\sqrt{2}\,dW_{t}^{i,N}+\sigma_{t}\,dB_{t},\quad$ (2)
-where $W_{t}^{i,N}$ and $B_{t}$ are independent standard $\mathbb{R}^{d}$-valued Brownian motions, defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$, $K$ is a singular kernel and $V^{N}$ a suitable scalling.
-To the best of the author’s knowledge, this is the first time that (1) is derived for 2D stochastic vortex model, by moderately interacting particles, in the whole space - extending what was known before for bounded kernels in [32], on $\mathbb{R}^{d}$, and in [39], for sub-Coulombian kernels, under periodic boundary conditions - with pathwise quantitative bounds, in the sense of the relative entropy functional.
-In the particle system (2), the stochastic term $\sigma_{t}dB_{t}$ accounts for a source of randomness that affects all particles, modeling a shared environmental fluctuation.
-Models of this nature appear in several contexts, such as problems in statistical mechanics, population dynamics and biology, as well as in the analysis of numerical Monte–Carlo schemes and related applications; see, e.g., [4], [5], [24], [19], and [29].
-The main goal of this paper is to investigate the asymptotics in $N$ of the empirical measure defined by
+where $W_{t}^{i,N}$ and $B_{t}$ are independent standard $\mathbb{R}^{d}$-valued Brownian motions, defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$, $K$ is a singular kernel, $V^{N}$ is a suitable scaling and $\sigma$ is a diffusion matrix.
+To achieve these bounds, we study the asymptotic behavior, as $N\to\infty$ of the empirical measure defined by
 $S_{t}^{N}\doteq\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},\qquad t\geqslant 0,$ (3)
-where $\delta_{a}$ is the delta Dirac measure concentrated at $a\in\mathbb{R}^{d}$.
-For that purpose, considering $V^{N}\doteq N^{\beta}V(N^{\frac{\beta}{d}}\cdot)$, we introduce the mollified empirical measure
+where $\delta_{a}$ is the Dirac delta measure concentrated at $a\in\mathbb{R}^{d}$, by introducing the mollified empirical measure
 $\rho_{t}^{N}\doteq V^{N}\ast S_{t}^{N}=\int_{\mathbb{R}^{d}}V^{N}(\cdot-y)\,S_{t}^{N}(dy),$
-in order to apply the relative entropy functional directly to the densities $\rho^{N}$ and $\rho$, since by the Csiszar-Kullback-Pinsker inequality, we get
-$\|S^{N}_{t}-\rho_{t}\|_{0}^{2}\lesssim N^{-\frac{\beta}{d}}+\|\rho^{N}_{t}-\rho_{t}\|_{1}^{2}\overset{\text{CKP}}{\lesssim}N^{-\frac{\beta}{d}}+\mathcal{H}\left(\rho^{N}_{t}|\rho_{t}\right),$ (4)
-where the metric in the left hand side of (4), is the Kantorovich-Rubinstein metric between two probatility measures.
-It follows that, if there exists a constant $C=C_{\omega}>0$ such that for any $N\in\mathbb{N}$,
-$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})\leq C(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}),$ (5)
-where $\theta$ is an explicit positive parameter, by (4) and (5), we have, in particular, propagation of chaos for the marginals of the empirical measure of the particle system, see [1, Section 8.3] and [45].
-In the present work, we apply for the first time in the context of moderately interacting particles, the Donsker–Varadhan inequality together with the dissipation provided by the Fisher information, to handle the nonlinearity.
-In addition, we employ localization techniques combined with the probabilistic data setting, to derive estimates for the quadratic variation terms in the whole space, as the estimate established in [39], under periodic boundary conditions, does not apply in the present framework.
+which allows us to apply the relative entropy functional directly to the densities $\rho^{N}$ and $\rho$.
+To the best of our knowledge, this is the first time that pathwise quantitative bounds, in the sense of the relative entropy functional, have been derived for the 2D stochastic vortex model in the whole space via moderately interacting particles.
+This extends previously known results for bounded kernels on $\mathbb{R}^{d}$ [33] and for sub-Coulombian kernels under periodic boundary conditions [40].
+As a consequence of our entropy bounds, we obtain energy estimates for the difference between the regularized empirical measure and the solution of (1), extending previous results by [27] and [37].
+Our main novelty lies in the key technical tools we develop.
+For the first time in the context of moderately interacting particles, we apply the Donsker–Varadhan inequality, together with the dissipation provided by the Fisher information, to handle the nonlinearity.
+In addition, we employ localization techniques combined with our probabilistic data setting to derive estimates for the quadratic variation terms in the whole space, since the estimates established in [40] under periodic boundary conditions do not apply in the present framework.
+Finally, new energy estimates quantifying the distance between the regularized empirical measure and the solution to (1) are derived by combining the control of the Fisher information of the particle system with the Ladyzhenskaya inequality and a nonlinear Grönwall inequality.
 Moreover, we prove the existence of a suitable solution to the equation (1).
@@
 The relative entropy method for establishing quantitative propagation of chaos in McKean–Vlasov systems with singular interaction kernels was first introduced in [24].
-It was developed in the context of general first-order systems with interaction kernels in $W^{-1,\infty}$, a framework that includes the point vortex model approximating the two-dimensional Navier–Stokes equation.
-We also refer to [42], where the mean-field limit and quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels are established via the modulated energy method.
+It was developed in the context of general first-order systems with interaction kernels in $W^{-1,\infty}$, a framework that includes the 2D viscous vortex model.
+We also refer to [43], where quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels are established via the modulated energy method.
 In contrast to the relative entropy approach, which operates at the level of the joint law of the particle system, the modulated energy method is formulated directly in terms of the empirical measure associated with the particles.
 Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results.
-In [2] and [3], the strategies given in [23] and [42], were adapted to get quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes.
+In [2] and [3], the strategies given in [24] and [43], were adapted to obtain a quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes.
 Additionally, [21], by means of a logarithmic Sobolev inequality, derived uniform in time propagation of chaos for the 2D vortex model, adapting the strategy in [24].
@@
 Regarding sub-Coulombian kernels, [6] obtained the derivation of the mean-field approximation for Landau-like equations, while [41] obtained global-in-time mean-field convergence for gradient (conservative) diffusive flows.
-By extending the BBGKY hierarchy given in [30], for singular kernels, [47] obtained the optimal convergence rate for the relative entropy between the k-particle marginals and the corresponding tensorized law for weakly interacting diffusions with bounded interaction kernels.
-With this in view we observe that, the relative entropy method has proven effective for deriving quantitative estimates, for singular kernels, under periodic boundary conditions, while the modulated energy method, which can be applied in the whole space, has primarily been developed for deterministic interaction.
-However, more recently, quantitative propagation of chaos for the 2D vortex model and the 2D Log Gas on the whole space was obtained in [48] and [49], [7], respectively.
+By extending the BBGKY hierarchy introduced in [31] to the case of singular kernels, and adapting the strategy developed in [24], [48] obtained the optimal convergence rate for the relative entropy between the k-particle marginals and the corresponding tensorized law.
+More recently, quantitative propagation of chaos for the 2D vortex model on the whole space was obtained in [49, 50], while the corresponding result for the 2D log-gas was established in [7].
 In these works, the authors exploit Li–Yau-type estimates and Hamilton-type heat kernel estimates to derive suitable bounds for the limiting solution and its derivatives up to second order, which allows them to extend the strategy given in [24], to $\mathbb{R}^{d}$.
-Quantitative estimates in the whole space were derived, in the context of moderately interacting particles  [^2] The moderate interact model was introduced by Oelschlager in [34],[33] and [35].
-with singular kernels - including the Biot-Savart, the Keller-Segel and the Dirac delta measure kernels - by [36] and [37], based on a semigroup approach introduced in [17], with further advancements in [8], [21], [38], and [27].
+Quantitative estimates in the whole space were derived, in the context of moderately interacting particles  [^2] The moderately interacting particle model was introduced by Oelschläger in [35, 34, 36].
+with singular kernels and without common noises - including the Biot-Savart, the Keller-Segel and the Dirac delta measure kernels - by [37] and [38], based on a semigroup approach introduced in [18], with further advancements in [8], [22], [28] and [39].
 It is important to note that these works do not rely on the relative entropy method, but instead establish quantitative estimates in strong functional topologies.
 The connection between the time evolution of the relative entropy, of the joint distribution, and moderate interacting regime, on the whole space was carried out by [10], based on a combination between the relative entropy and the regularised $L^{2}$-estimate in [36], deriving a propagation of chaos result for the viscous porous medium equation.
-But was in [39], on the torus, that the equation (1), was derived, for singular kernels and common noise, in the context of moderate interaction regime, by applying the relative entropy directly to the mollified empirical measure.
-Earlier, also the same domain, the results established in [36] were extended to the setting with environmental noise in [26].
-It is worth highlighting, that quantitative entropy estimates for systems with individual and common noise, as shown in [43] for incompressible the Navier-Stokes equations, [9] for the Hegselmann-Krause model, and [32] for mean-field systems with bounded kernels.
+However, it was in [40] that equation (1) was derived on the torus for singular kernels and common noise in the moderately interacting regime, by applying the relative entropy directly to the mollified empirical measure.
+Earlier, also in the same domain, the results established in [37] were extended to the setting with environmental noise in [27].
+It is worth highlighting, that quantitative entropy estimates for systems with individual and common noise, have been obtained, as shown in [44] for incompressible the Navier-Stokes equations, [9] for the Hegselmann-Krause model, and [33] for mean-field systems with bounded kernels.
 For more results, see [19], [15], [16], and [29].
+Particle systems driven by environmental noises, such as given by (2), appear in several contexts, such as problems in statistical mechanics, population dynamics and biology, as well as in the analysis of numerical Monte–Carlo schemes and related applications; see, e.g., [4], [5], [25], [20], and [30].
 In contrast to the aforementioned works, our application of the relative entropy method provides pathwise estimates, deriving bounds directly at the trajectory level of the particle system.
 In those works, the estimates are obtained at the level of the joint law.
+Outline of the article
 In order to present our results, this work is organized as follows.
@@
 In the subsection 1.3, we present our main results.
-The section 2 is dedicated to the proof of our main results.
-To ease the presentation, the Appendix, provides the corresponding suitable notion of a solution for the limiting equation, as well as some inequalities used throughout the proofs.
+The sections 2 and 3 are dedicated to the proofs of our main results.
+To ease the presentation, the Appendix, provides the existence of a suitable solution for the limiting equation, as well as some inequalities used throughout the proofs.
 1.1 Notations
@@
 For a normed vector space $U$, we denote its norm by $\|\cdot\|_{U}$, except for $L^{p}$ spaces, where we write $\|\cdot\|_{p}$, for $p\geq 1$.
+For $\gamma\in(0,1]$ the Hölder space on $\mathbb{R}^{d}$ is given by
+$C^{\gamma}=C^{\gamma}(\mathbb{R}^{d})\doteq\left\{f:\mathbb{R}^{d}\to\mathbb{R}^{e}\mid\|f\|_{\gamma}\doteq\|f\|_{\infty}+\sup_{x,y\in\mathbb{R}^{d}}\frac{|f(x)-f(y)|}{|x-y|^{\gamma}}<\infty\right\}.$
 The Kantorovich-Rubinstein metric reads, for any two probability measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$,
@@
   1. $(\mathbf{A}^{V})$
-    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)=N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V:\mathbb{R}^{d}\to\mathbb{R}$ satisfying
+    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)=N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V\in\mathcal{P}\left(\mathbb{R}^{d}\right)$ satisfying
     $\left|\nabla V\right|\leq C_{d}V,$ (5)
-    and for $\alpha>d/2$
-    $V(y)\leq C_{d}(1+|y|^{2})^{-\alpha}.$ (8)
+    for $\alpha>d/2$
+    $V(y)\leq C_{d}(1+|y|^{2})^{-\alpha},$ (6)
+    and for $\gamma\in(0,1)$ and $y\neq 0$
+    $V(y)\leq C_{d}|y|^{-3d-\gamma}.$ (7)
+  2. $(\mathbf{A}^{K})$
+    It holds for $\gamma\in(1/2,1)$ that
+    $\|K\ast f\|_{\gamma}\leq C\left(\|f\|_{1}+\|f\|_{4}\right)$ (8)
+    for all $f\in H^{1}$.
   3. $(\mathbf{A}^{\nabla\cdot K})$
@@
   5. $(\mathbf{A}^{\rho_{0}})$
-    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,\infty}\left(\mathbb{R}^{d}\right)$ and
+    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,\infty}\cap\mathcal{P}\left(\mathbb{R}^{d}\right)$ and
     $\left|\nabla\ln{\rho_{0}}(x)\right|^{2}\leq\widetilde{C}_{1}(1+|x|^{2}),$ (10)
@@
   6. $(\mathbf{A}^{X_{0}})$
-    The initial particle system $(X_{0}^{i,N})_{i=1}^{N}$, $N\in\mathbb{N}$ is a sequence of i.i.d with law $\rho_{0}$.
+    The initial particle system $(X_{0}^{i,N})_{i=1}^{N}$, $N\in\mathbb{N}$ is a sequence of i.i.d random variables with law $\rho_{0}$.
 Remark 1.
-For an example of mollifier $V:\mathbb{R}^{d}\to\mathbb{R}$ satisfying $\left(\mathbf{A}^{V}\right)$, take
+For an example of mollifier $V\in\mathcal{P}(\mathbb{R}^{d})$ satisfying $\left(\mathbf{A}^{V}\right)$, take
 $V(y)\doteq\overline{C}_{d}\exp{\left(-\sqrt{1+|y|^{2}}\right)}$
@@
 $K(x)=\frac{1}{2\pi}\frac{x^{\perp}}{|x|^{2}}$ (13)
-and then, it is divergence in the sense of distributions, of a $L^{\infty}$ matrix $K_{0}$, namely $K=\nabla\cdot K_{0}$, with
-$K_{0}(x_{1},x_{2})=\left(\frac{1}{2\pi}\arctan{\left(\frac{x_{1}}{x_{2}}\right)}\right)I.$ (14)
-It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$.
+and then, it is the divergence in the sense of distributions, of an $L^{\infty}$ matrix $K_{0}$, namely $K=\nabla\cdot K_{0}$, with
+$K_{0}(x_{1},x_{2})=-\left(\frac{1}{2\pi}\arctan{\left(\frac{x_{1}}{x_{2}}\right)}\right)I.$ (14)
+It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$. Furthermore, $K$ satisfy $(\mathbf{A}^{K})$, see [39].
 Remark 3.
@@
 1.3 Statement of the main results
-The first result, whose proof is postponed to the Appendix, establishes that the limiting equation is well posed in the appropriate functional framework.
-Indeed, it provides the regularity properties required in the proof of the main results.
-Theorem 1.
-Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. Then, there exist $T_{1}>0$, $C_{3}>0$, such that if $\widetilde{C}_{3}<C_{3}$,
-  • there exists a pathwise solution of (1) with initial condition $\rho_{0}$, such that $\rho\in L^{\infty}\left([0,T_{1}];W^{2,1}\cap W^{2,\infty}(\mathbb{R}^{d})\right)$, $\mathbb{P}$-a.s.;
+In this subsection, we establish our main results. We begin by formulating a suitable notion of solution, needed for deriving quantitative entropy estimates.
+Definition 1.1.
+Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$.
+A pathwise solution for (1) with initial condition $\rho_{0}$, is a measurable function $\rho:[0,T_{1}]\times\mathbb{R}^{d}\times\Omega\to\mathbb{R}$, for some $T_{1}>0$, such that:
+  • it holds $\mathbb{P}$-a.s.
+    $\rho\in L^{\infty}\left([0,T_{1}];W^{2,1}\cap W^{2,\infty}\cap\mathcal{P}(\mathbb{R}^{d})\right)$ (15)
   • the solution $\rho$ verifies, $\mathbb{P}$-a.s., that
@@
     and
-    $\rho(t,x)\leq 2\exp{(-2\pi|x-X_{t}|^{2})},$ (17)
+    $\rho(t,x)\leq\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)},$ (18)
     where
     $dX_{t}=\sigma_{t}dB_{t},$ (19)
-    for all $t\in[0,T_{1}]$ and constants $C_{1},C_{2}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
+    for all $t\in[0,T_{1}]$ and constants $C_{1},C_{2}$ and $C_{3}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
 We are now able to extend what was known before for bounded kernels in the whole space as in [33] and in [40], for sub-Coulombian kernels, under periodic boundary conditions.
@@
 Theorem 1.
-Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$ and let $\rho$ be a solution of (1), given by Theorem 1.
+Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$ and let $\rho$ be a solution of (1), given by Definition 1.1.
 In addition, let the dynamics of the particle system be given by (2) and
-$\mathcal{H}(\rho_{0}^{N}|\rho_{0})\leq C,\,\,\,\mathbb{P}-a.s.$
-where
+$\lim_{N\to\infty}N^{\theta}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.$ (20)
+and
 $\theta=\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
 with $\delta>0$, such that $\theta>0$, $d\geq 1$ and $\beta\in\left(0,\frac{1}{2\big[1+\frac{1}{d}\big]}\right)\cap\left(0,\frac{1}{\big[1+\frac{2}{d}+2\alpha\big]}\right)$.
-Then, there exist a time $T\in(0,T_{1})$ and a constant $C>0$, such that, almost surely for $N>>1$,
-$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (19)
-As a corollary of the previous theorem, we establish propagation of chaos for the marginals of the empirical measure of the particle system.
-In the sense of [1, Section 8.3], it corresponds to convergence of the empirical measure with respect to the Kantorovich–Rubinstein metric, given in (6).
+Then, there exist a time $T\in(0,T_{1})$, an almost surely finite random variable $A_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
+$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (21)
+In particular, we get
+$\limsup_{N\to\infty}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim T,\,\,\,\,\,\,\mathbb{P}-a.s.$ (22)
+The following result establishes energy estimates as an application of our entropy bounds, extending previous results by [27] and [37].
+Precisely, we utilize the entropy dissipation to control the Fisher information of the particle system, which in turn allows us to derive, a new energy estimate for the difference between the regularized empirical measure and the solution to (1), using the Ladyzhenskaya inequality and a nonlinear Gronwall lemma.
+Theorem 2.
+Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{K})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$, $d=2$ and let $\rho$ be a solution of (1), given by Definition 1.1.
+In addition, let the dynamics of the particle system be given by (2) and
+$\lim_{N\to\infty}N^{\widetilde{\theta}}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.,$ (23)
+$\lim_{N\to\infty}N^{\widetilde{\theta}}\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}=0,\,\,\,\mathbb{P}-a.s.$ (24)
+and
+$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
+where
+$\theta_{1}\doteq\left(1-\beta\left(2+2\alpha\right)\right);\,\,\,\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right);\,\,\,\widetilde{\theta}_{1}\doteq\left(1-2\beta(1+\frac{1}{2})+\beta\right)$
+$\widetilde{\theta}_{2}\doteq\beta(2-3\gamma);\,\,\,\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{2};\,\,\,\widetilde{\theta}_{4}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right)$
+with $\gamma\in(1/2,2/3)$ and $\delta>0$, such that $\widetilde{\theta}>0$ and $\beta\in\left(0,\frac{1}{2+2\alpha}\right)$.
+Then, there exist a time $T\in(0,T_{1})$, a pair of almost surely finite random variables $A_{0}$, $\widetilde{A}_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
+$\sup_{t\in[0,T]}\|\rho_{t}^{N}-\rho_{t}\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
+[…]
+In particular, we get
+$\limsup_{N\to\infty}\left(\sup_{t\in[0,T]}\|\rho_{t}^{N}-\rho_{t}\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds\right)$ $\lesssim T,\,\,\,\mathbb{P}-a.s.$ (26)
+As a corollary of Theorem 1, we establish quantitative bounds for the marginals of the empirical measure of the particle system.
+In the sense of [1, Section 8.3], it corresponds to get bounds of the empirical measure in the sense to the Kantorovich–Rubinstein metric, given in (4).
 Corollary 1.
-Under assumptions of Theorem 2, we have
-$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (20)
+Under the assumptions of Theorem 1, for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$, such that for $N\geq N_{0}$ we have
+$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (27)
+In particular, we get
+$\limsup_{N\to\infty}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim T,\,\,\,\mathbb{P}-a.s.$ (28)
 Remark 4.
 We observe that, compared with previous works on moderate particle systems without common noise, we lose the order of convergence.
-This is because the estimation of the martingale term in our work is of order $\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)$, while in [36] is the order $\frac{1}{2}-\frac{\beta}{2}$, see Theorem 1.3 in [36].
+This is because the estimation of the martingale term in our work is is of order $\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)$, while in [37] is the order $\frac{1}{2}-\frac{\beta}{2}$, see Theorem 1.3 in [37].
 When environmental noise is taken into account, the overall framework remains comparable to that of [27].
 For the same reason, the range of the parameter $\beta$ is smaller in our work.
-However, in the present work we obtain an improved convergence rate compared to that established in [39].
-2 Proofs of main results
-In this section we present the proofs of our main results.
+However, in the Theorem 1, we obtain an improved convergence rate compared to that established in [40].
+The next result, whose proof is postponed to the Appendix, establishes that the limiting equation is well posed in the appropriate functional framework, given by Definition 1.1.
+Theorem 3.
+Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. Then, there exists a solution of (1) in the sense of Definition 1.1.
+2 Proofs of Theorem 1
+In this section we present the proofs of Theorem 1.
 In Subsection 2.1, by applying the Itô’s formula, such as in [40], we derive an evolution equation for the relative entropy functional of the regularized empirical measure with respect to the solution of (1).
@@
 $\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
-$\overset{\alpha>d/2}{\leq}C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}.$ (38)
+$\overset{\alpha>d/2}{\leq}C_{\alpha,T}N^{-\theta_{1}},$ (46)
+for $\theta_{1}\doteq\left(1-\beta\left(1+2/d+2\alpha\right)\right)$.
 2.3 Estimates for $I_{t}$: the nonlinear terms
 We now address the nonlinear terms in (29).
-The main novelty is to use Donsker-Varadhan inequality (82), in the context of moderate interactions, along with the dissipation that comes from of Fisher information and Assumption $(\mathbf{A}^{\nabla\cdot K})$, to tackle with the singularity of kernel $K$.
+The main novelty is to use Donsker-Varadhan inequality (108), combined with the notion of solution of limiting process given by Definition 1.1, in the context of moderate interactions, along with the dissipation that comes from of Fisher information and Assumption $(\mathbf{A}^{\nabla\cdot K})$, to tackle with the singularity of kernel $K$.
 First, we set
 $I_{t}$ $\overset{}{=}\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho_{s}^{N}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)[K\ast\rho_{s}-K\ast\rho_{s}^{N}]\,dx\,ds$
-$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
+$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\underbrace{\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle}_{\doteq I^{N}}\,dx\,ds$
 $\doteq I^{1}_{t}-I^{2}_{t},$ (47)
@@
 $\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\left|K_{0}\ast(\rho_{s}-\rho_{s}^{N})\right|^{2}\,dx\,ds$
-$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\|K_{0}\|\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\,dx\,ds.$ (41)
-We observe that by Donsker-Varadhan inequality (82), for $\eta>>1$
+$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\|K_{0}\|_{\infty}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\,dx\,ds.$ (49)
+We observe that, for $\eta>2C_{1}(8T+C_{3})+1$
+$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty$ (50)
+which implies by Donsker-Varadhan inequality (108),
 […]
 […]
-$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
-$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$
-which implies by Csiszár-Kullback-Pinsker inequality (81) and (41)
+Hence by Csiszár-Kullback-Pinsker inequality (107) and (49), we derive
 $I_{t}^{1,2}$ $\leq\epsilon\int_{0}^{t}\mathcal{I}(\rho^{N}_{s}|\rho_{s})\,ds+C_{\epsilon,\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds.$ (52)
@@
 […]
-by Donsker-Varadhan inequality (82), we obtain
+for $\eta>2C_{1,2}(8T+C_{3})+1$, we obtain
+$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty.$ (55)
+It follows by Donsker-Varadhan inequality (108),
 […]
 […]
-$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
-$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$ (45)
-for $\eta>>1$.
 Therefore, by (56) in (53), we get
@@
 $\overset{}{=}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}.$ (60)
-Now from Leibniz’s rule we deduce, for $\eta>>1$
+Now from Leibniz’s rule we deduce,
 $\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx$ $=\int_{\mathbb{R}^{d}}\frac{\left|\nabla\left(\rho_{s}\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
@@
 $\overset{}{\leq}2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\frac{\left|\nabla\rho_{s}\right|^{2}}{\rho_{s}^{2}}\,dx+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx$
-$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (50)
-where in the last inequality, we are using Donsker-Varadhan inequality (82):
 […]
 […]
-$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
-$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}.$
-[…]
-$\left|\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\right|$ $\leq 2\|K_{0}\|_{\infty}\left\langle S_{s}^{N},V^{N}(x-\cdot)\right\rangle$
-$\times\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}$
+$I^{N}$ $\leq 2\|K_{0}\|_{\infty}\left\langle S_{s}^{N},V^{N}(x-\cdot)\right\rangle\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}$
 $\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}},$
@@
 $+\sqrt{2}/2\int_{0}^{t}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\,dx\,ds$
-$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\eta}$
-$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\eta}.$ (51)
+$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\epsilon,\eta}$
+$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\epsilon,\eta}.$ (62)
 […]
-$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (52)
+$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\epsilon,\eta}.$ (63)
 2.4 Application of Gronwall’s Lemma
-In this subsection, we conclude our estimates, by applying Gronwall’s inequality.
-Indeed, with the same computations as in [39], for all $\delta\in(0,\theta_{1})$ there exists a random variable $A_{0}$, finite almost surely, such that
-$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{1}+\delta},$
-with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
+In this subsection, we conclude our localizated estimates, by applying Gronwall’s inequality.
+Indeed, with the same computations as in [40], for all $\delta\in(0,\theta_{2})$ there exists a random variable $A_{0}$, finite almost surely, such that
+$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{2}+\delta},$ (64)
+with $\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
 […]
@@
 $-\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
-$+C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}$
-$+A_{0}N^{-\theta_{1}+\delta}$
-$+TC_{\eta}$ (53)
+$+C_{\alpha,T}N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+TC_{\eta}$ (65)
 and then for $\eta>>1$ and $\epsilon<<1$, we deduce
-$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{2}}+A_{0}N^{-\theta_{1}+\delta}+C,$ (54)
-with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$ and $\theta_{2}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$.
-Finally, from Gronwall’s Lemma in (54), recalling (28), (31) and (32), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
-$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (55)
+$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+T,$ (66)
+with $\theta_{1}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$ and $\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
+Finally, from Gronwall’s inequality (109) in (66), recalling (36), (39) and (40), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
+$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right)$ (67)
 with
@@
 We now remove the stopping time $\tau^{N}$ introduced in (67) by exploiting the probabilistic framework provided by Assumptions $\left(\mathbf{A}^{\rho_{0}}\right)$ and $\left(\mathbf{A}^{X_{0}}\right)$.
-These assumptions ensure sufficient control on the initial data and on the particle system, allowing us to pass from the localized estimates to global-in-time bounds with high probability, and finally eliminate the stopping time.
+It allows us to pass from the localized estimates to uniform bounds with high probability, removing the stopping time.
 First, since $K=\nabla\cdot K_{0}$, $K_{0}\in L^{\infty}$ and $V^{N}=N^{\beta}V(N^{\beta/d}\cdot)$, by (2), we have
@@
 Finally, by (40) and (80), for $T<\min{\left(T_{1};(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})\right)}$ for all $\omega\in\Lambda$, with
-$\Lambda\doteq\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}\right)\cap\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}\{\tau^{N}=T\}\right),$ (69)
-there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)=T$ and (55) holds. It follows that, almost surely
-$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (70)
-which conclude the proof of Theorem.
+$\Lambda\doteq\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}\right)\cap\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}\{\tau^{N}=T\}\right)\cap\left\{A_{0}<\infty\right\},$ (81)
+there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, we have $\tau^{N}(\omega)=T$ and
+$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (82)
+In particular, by (20) and (64), we conclude with
+$\limsup_{N\to\infty}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim T,$ (83)
+almost surely.
+3 Proof of Theorem 2
+In this section we present the proofs of Theorem 2.
+In Subsection 3.1, by applying the Itô’s formula, such as in [27], we derive an evolution equation for the energy of the difference between regularized empirical measure and the solution of (1).
+In Subsection 3.2, we address the nonlinear terms, combining bounds on the Fisher information of the particle system with the Ladyzhenskaya inequality.
+Finally, in Subsections 3.3, combining the previous estimates with a nonlinear Gronwall’s lemma, we complete the proof.
+3.1 Time evolution of energy
+Applying Itô’s lemma, following the approach of [27], we obtain
+$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}$ $\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}+\widetilde{I}^{1}_{t}-\widetilde{M}_{t}^{N}-2\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
+$+TN^{-\widetilde{\theta}_{1}}\,\|\nabla V\|_{2}^{2},$ (84)
+with $\widetilde{\theta}_{1}\doteq\left(1-2\beta(1+\frac{1}{d})+\beta\right)$,
+$\widetilde{I}^{1}_{t}\doteq 2\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla(\rho_{s}$ $-\rho_{s}^{N})\Big[\rho_{s}K\ast\rho_{s}-\left\langle S_{s}^{N},V^{N}(x-\cdot)K\ast\rho^{N}_{s}(\cdot)\right\rangle\Big]\,dx\,ds,$
+and
+$\widetilde{M}^{N}_{t}\doteq\frac{2\sqrt{2}}{N}\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{R}^{d}}(\rho_{s}-\rho_{s}^{N})\,(\nabla V^{N})(x-X_{s}^{i,N})\,dx\,dW_{s}^{i,N}.$
+3.2 Estimates for $\widetilde{I}_{t}$: the nonlinear terms
+Now by by subtracting and adding the term $\rho_{s}^{N}K\ast\rho_{s}^{N}$, along with integration by parts, we have
+$2\int_{\mathbb{R}^{2}}(\rho_{s}-\rho_{s}^{N})\left[\nabla\cdot\left[\rho_{s}K\ast\rho_{s}\right]-\nabla\cdot\left[\left\langle S_{s}^{N},V^{N}(x-\cdot)K\ast\rho_{s}^{N}(\cdot)\right\rangle\right]\right]\,dx$
+$=-2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left[\rho_{s}K\ast\rho_{s}-\left\langle S_{s}^{N},V^{N}(x-\cdot)K\ast\rho_{s}^{N}(\cdot)\right\rangle\right]\,dx$
+$=-2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left[\rho_{s}K\ast\rho_{s}-\rho_{s}^{N}K\ast\rho^{N}_{s}\right]\,dx$
+$+2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left\langle S_{s}^{N},V^{N}(x-\cdot)\left[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right]\right\rangle\,dx$
+$\doteq\widetilde{I}^{1,1}+\widetilde{I}^{1,2}.$ (85)
+In addition, since $\nabla\cdot K=0$, by by subtracting and adding the terms $\rho_{s}$ and $\rho_{s}^{N}K\ast\rho_{s}$, we find
+$\left|\widetilde{I}^{1,1}\right|$ $\leq\left|\underbrace{2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})(\rho_{s}-\rho_{s}^{N})K\ast\rho_{s}\,dx}_{=0}\right|$
+$+\left|\underbrace{2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})(\rho_{s}-\rho_{s}^{N})K\ast(\rho_{s}-\rho_{s}^{N})\,dx}_{=0}\right|$
+$+\left|2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\rho_{s}K\ast(\rho_{s}-\rho_{s}^{N})\,dx\right|$
+[…]
+By Young’s inequality, we observe that
+$\|\rho_{s}-\rho_{s}^{N}\|_{1}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}\leq C_{\epsilon}\|\rho_{s}\|_{2}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (87)
+By Ladyzhenskaya and Young inequalities, we deduce
+$\|\rho_{s}-\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$ $\lesssim\|\rho_{s}-\rho_{s}^{N}\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
+$\leq C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (88)
+Therefore by (87) and (88) in (86), we get
+$\widetilde{I}^{1,1}\leq C_{\epsilon}\|\rho_{s}\|_{2}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+2\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (89)
+Now we will take care about $\widetilde{I}^{1,2}$ in (85). First, taking into account (7), with analogous computations such as in section 2 of [40], we derive
+$\left|\left\langle S^{N}_{s},V^{N}(x-\cdot)|x-\cdot|^{\gamma}\right\rangle\right|\leq C_{d}N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}(x)$ (90)
+with $\widetilde{\theta}_{2}\doteq\beta(2-3\gamma)$ and $\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{d}$, for $\gamma\in(1/2,2/3)$.
+It follows by (8), that
+[…]
+[…]
+$\lesssim 2\|\rho_{s}^{N}\|_{1}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)$
+$+2\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)$
+$\doteq\widetilde{I}^{1,2,1}+\widetilde{I}^{1,2,2}$ (91)
+Now, by by subtracting and adding $\rho_{s}$, we note that
+$\widetilde{I}^{1,2,1}$ $\leq N^{-\widetilde{\theta}_{2}}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)$
+$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}-\rho_{s}^{N}|$
+$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}|$
+$\overset{\text{Young}}{\leq}N^{-\widetilde{\theta}_{2}}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}+C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{2}\right).$ (92)
+Again, by by subtracting and adding $\rho_{s}$, we find
+$\widetilde{I}^{1,2,2}$ $\leq N^{-\widetilde{\theta}_{2}}\|\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}+N^{-\widetilde{\theta}_{3}}\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\rho_{s}^{N}$
+$\overset{\text{Holder}}{\leq}N^{-\widetilde{\theta}_{2}}\|\rho_{s}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}+N^{-\widetilde{\theta}_{2}}\|\rho_{s}^{N}-\rho_{s}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}$
+$+N^{-\widetilde{\theta}_{3}}\|\rho_{s}-\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}-\rho_{s}^{N}\|_{2}$
+$+N^{-\widetilde{\theta}_{3}}\|\rho_{s}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}-\rho_{s}^{N}\|_{2}$
+$+N^{-\widetilde{\theta}_{3}}\|\rho_{s}-\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
+$+N^{-\widetilde{\theta}_{3}}\|\rho_{s}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
+$\doteq\widetilde{I}^{1,2,2}_{1}+\widetilde{I}^{1,2,2}_{2}+\widetilde{I}^{1,2,2}_{3}+\widetilde{I}^{1,2,2}_{4}+\widetilde{I}^{1,2,2}_{5}.$ (93)
+Now by Ladyzhenskaya and Young inequalities, we find
+$\widetilde{I}^{1,2,2}_{1}$ $\overset{\text{Young}}{\leq}N^{-\widetilde{\theta}_{2}}\left(\|\rho\|_{4}^{2}+\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}^{2}+\|\rho_{s}-\rho_{s}^{N}\|_{4}^{2}\right)$
+$\overset{\text{Young}+\text{Lady}}{\leq}N^{-\widetilde{\theta}_{2}}\left(\|\rho_{s}\|_{4}^{2}+\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}^{2}\right)$
+$+N^{-\widetilde{\theta}_{2}}\left(C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}\right).$ (94)
+Analogously we deduce
+$\widetilde{I}^{1,2,2}_{3}+\widetilde{I}^{1,2,2}_{4}+\widetilde{I}^{1,2,2}_{5}$ $\leq N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{4}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{4}^{4}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\right).$ (95)
+Also, for the term $\widetilde{I}^{1,2,2}_{2}$, by means of Ladyzhenskaya and Young inequalities we get
+$\widetilde{I}^{1,2,2}_{2}\leq N^{-\widetilde{\theta}_{3}}\left(C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{6}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}\right).$ (96)
+It follows by (92), (93), (94), (95) and (96) in (91), we arrive at
+$\widetilde{I}^{1,2}$ $\overset{}{\leq}N^{-\widetilde{\theta}_{2}}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}+C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{2}\right)$
+$\overset{}{+}N^{-\widetilde{\theta}_{2}}\left(\|\rho\|_{4}^{2}+\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}^{2}\right)$
+$+N^{-\widetilde{\theta}_{2}}\left(C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{4}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{4}^{4}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{6}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}\right).$ (97)
+Hence, by (89) and (97) in (85), we conclude with
+$\widetilde{I}_{t}$ $\leq C_{\epsilon}\|\rho_{s}\|_{2}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+2\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}$
+$\overset{}{+}N^{-\widetilde{\theta}_{2}}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}+C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{2}\right)$
+$\overset{}{+}N^{-\widetilde{\theta}_{2}}\left(\|\rho\|_{4}^{2}+\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}^{2}\right)$
+$+N^{-\widetilde{\theta}_{2}}\left(C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{4}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+C_{\epsilon}\|\rho_{s}\|_{4}^{4}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\right)$
+$+N^{-\widetilde{\theta}_{3}}\left(C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{6}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}\right).$ (98)
+3.3 End of the proof: application of Gronwall Lemma
+First, with the same computations as in [40], for all $\delta\in(0,\widetilde{\theta}_{4})$ there exists a random variable $\widetilde{A}_{0}$, finite almost surely, such that
+$\sup_{t\in[0,T]}|\widetilde{M}_{t}^{N}|\leq\widetilde{A}_{0}N^{-\widetilde{\theta}_{4}+\delta},$ (99)
+with $\widetilde{\theta}_{4}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
+Then by (98) and (99) in (84), for $\epsilon<<1$, we obtain
+$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$ $\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}$
+$+C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\,ds$
+$+\int_{0}^{t}\left(C_{\epsilon}+\|\rho_{s}\|_{4}^{2}+\|\rho_{s}\|_{2}^{4}\right)\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\,ds$
+$+C_{\epsilon}\int_{0}^{t}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{6}\,ds$
+$\overset{}{+}N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)\,ds$
+$\overset{}{+}2N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(\|\nabla\rho_{s}\|_{1}^{2}+\|\nabla\rho_{s}^{N}\|_{1}^{2}\right)\,ds$
+$+N^{-\widetilde{\theta}_{3}}C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\,ds+N^{-\widetilde{\theta}_{1}}\int_{0}^{t}\|\rho_{s}\|_{4}^{2}\,ds$
+$+N^{-\widetilde{\theta}_{3}}\int_{0}^{t}\left(C_{\epsilon}\|\rho_{s}\|_{4}^{4}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\right)\,ds$
+$+TN^{-\widetilde{\theta}_{1}}\,\|\nabla V\|_{2}^{2}$
+$+\widetilde{A}_{0}N^{-\widetilde{\theta}_{4}+\delta}.$ (100)
+Now by (18) $\|\rho\|_{T,2},\|\rho\|_{T,4}\leq C_{0}$, almost surely, and thus for a universal constant $C>0$ in (100) along side (60), (2.3) and Csiszár-Kullback-Pinsker inequality (107), we derive
+$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
+$\hskip 100.37498pt\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}$
+$\hskip 100.37498pt+\int_{0}^{t}C\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\int_{0}^{t}C\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\,ds$
+$\hskip 100.37498pt+\int_{0}^{t}C\|\rho_{s}-\rho_{s}^{N}\|_{2}^{6}\,ds$
+$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\nabla\rho_{s}\|_{1}\,ds+CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\nabla\rho_{s}\|_{1}^{2}\,ds$
+$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{1}}$
+$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+1+2\mathcal{I}(\rho_{s}^{N}|\rho_{s})\right)\,ds$
+$\hskip 100.37498pt+CN^{-\widetilde{\theta}_{3}}+CN^{-\widetilde{\theta}_{2}}$
+$\hskip 100.37498pt+CN^{-\widetilde{\theta}_{1}}\,\|\nabla V\|_{2}^{2}$
+$\hskip 100.37498pt+C\widetilde{A}_{0}N^{-\widetilde{\theta}_{4}+\delta}.$ (101)
+Finally, by adding (65) and (101), we deduce for $\epsilon<<1$ and $N>\left(\frac{4}{\epsilon}\right)^{1/\widetilde{\theta}_{1}}$
+$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})+\left\|\rho_{t\wedge\tau^{N}}-\rho^{N}_{t\wedge\tau^{N}}\right\|_{2}^{2}+\int_{0}^{t\wedge\tau^{N}}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
+$\leq\mathcal{H}(\rho_{0}^{N}|\rho_{0})+\|\rho_{0}-\rho_{0}^{N}\|_{2}^{2}+C_{N}+TC$
+$+\int_{0}^{t\wedge\tau^{N}}C\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\left(\int_{0}^{s}\left\|\nabla(\rho_{r}-\rho^{N}_{r})\right\|_{2}^{2}\,dr\right)\right)\,ds$
+$+\int_{0}^{t\wedge\tau^{N}}C\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\left(\int_{0}^{s}\left\|\nabla(\rho_{r}-\rho^{N}_{r})\right\|_{2}^{2}\,dr\right)\right)^{3}\,ds,$ (102)
+with
+$C_{N}\doteq CN^{-\widetilde{\theta}}\|\nabla\rho\|_{T,1}+CN^{-\widetilde{\theta}}\|\nabla\rho\|_{T,1}^{2}+CN^{-\widetilde{\theta}}+CA_{0}N^{-\widetilde{\theta}}+C\widetilde{A}_{0}N^{-\widetilde{\theta}}$,
+and
+$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
+for $\delta<<1$.
+It follows by (15), (23), (24) and (81), (99) in (102), that there exist $T>0$ and $\widetilde{\Lambda}\in\mathcal{F}$, $\mathbb{P}(\widetilde{\Lambda})=1$, with the following property: for all $\omega\in\widetilde{\Lambda}$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have $N^{-\widetilde{\theta}}+C_{N}+TC<\exp{(-CT)}(2CT)^{-1/2}$ and
+$\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
+$\hskip 20.075pt\leq N^{-\widetilde{\theta}}+C_{N}+TC$
+$\hskip 20.075pt+\int_{0}^{t}C\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\left(\int_{0}^{s}\left\|\nabla(\rho_{r}-\rho^{N}_{r})\right\|_{2}^{2}\,dr\right)\right)\,ds$
+$\hskip 20.075pt+\int_{0}^{t}C\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\left(\int_{0}^{s}\left\|\nabla(\rho_{r}-\rho^{N}_{r})\right\|_{2}^{2}\,dr\right)\right)^{3}\,ds.$ (103)
+In particular by the nonlinear Gronwall’s inequality (111), for all $t\in[0,T]$, we deduce
+$\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$ $\leq N^{-\widetilde{\theta}}+C_{N}+TC$ (104)
+and thus, we end up with
+$\limsup_{N\to\infty}\sup_{t\in[0,T]}\left(\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds\right)$ $\lesssim T,$ (105)
+almost surely.
 4 Appendix
 This appendix contains the proof of Theorem 1 together with several auxiliary inequalities used in the previous sections.
-We begin by establishing Theorem 1, and subsequently derive the additional estimates required throughout the paper.
+We begin by establishing Theorem 1, regarding the existence of a suitable solution of the limiting equation.
 4.1 Proof of Theorem 3
-3.1.1 Solving (1) for $\sigma=0$
-When $\sigma=0$, a solution for (1) in the sense of Theorem 1, with exception of (17), was constructed in Sections 3 and 4 of [48] and [49].
-With this in view, we only need to verify that, it is possible to take $\tilde{C}_{3}$ sufficiently small, in order to get that Gaussian decay.
-Proof of (17).
-By Theorem 3 in [13] along with (12), we obtain there exists $C>0$ such that for all $(t,x)\in[0,T_{1}]\times\mathbb{R}^{d}$
-$\bar{\rho}(t,x)\leq C\widetilde{C}_{3}\int_{\mathbb{R}^{d}}\frac{1}{t^{d/2}}\exp{\left(-\frac{\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}}{8t\widetilde{C}_{3}}\right)}dy.$ (71)
-Now since
-$|x-y|^{2}=|x|^{2}-2x\cdot y+|y|^{2}.$
-we find
-$\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}$ $=\widetilde{C}_{3}|x|^{2}-2\widetilde{C}_{3}x\cdot y+\widetilde{C}_{3}|y|^{2}+8t|y|^{2}$
-$=\widetilde{C}_{3}|x|^{2}-2\widetilde{C}_{3}x\cdot y+(\widetilde{C}_{3}+8t)|y|^{2}.$ (72)
-Additionally, we get
-$(\widetilde{C}_{3}+8t)|y|^{2}-2\widetilde{C}_{3}x\cdot y.=(\widetilde{C}_{3}+8t)\left(|y|^{2}-2\frac{\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}x\cdot y\right).$
-and since
-$|y|^{2}-2ax\cdot y=|y-ax|^{2}-a^{2}|x|^{2},$
-with
-$a=\frac{\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}.$
-we arrive at
-$\Big((\widetilde{C}_{3}+8t)|y|^{2}-2\widetilde{C}_{3}x\cdot y\Big)=(\widetilde{C}_{3}+8t)|y-ax|^{2}-\frac{\widetilde{C}_{3}^{2}}{\widetilde{C}_{3}+8t}|x|^{2}.$ (73)
-Therefore by (72) and (73) we derive
-$\exp\left(-\frac{\widetilde{C}_{3}|x-y|^{2}+8t|y|^{2}}{8t\widetilde{C}_{3}}\right)=\exp{\left(-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}\right)}\exp\left(-\frac{\widetilde{C}_{3}+8t}{8t\widetilde{C}_{3}}|y-ax|^{2}\right).$ (74)
-and letting z=yaxz=y-ax we have
-$\bar{\rho}(x,t)$ $\leq C\widetilde{C}_{3}\frac{1}{t^{d/2}}e^{-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}}\int_{\mathbb{R}^{d}}\exp{\left(-\frac{\widetilde{C}_{3}+8t}{8t\widetilde{C}_{3}}|z|^{2}\right)}dz$
-$\leq C\widetilde{C}_{3}\frac{1}{t^{d/2}}e^{-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}}\left(\frac{\pi 8t\widetilde{C}_{3}}{\widetilde{C}_{3}+8t}\right)^{d/2}$
-$\leq C\widetilde{C}_{3}(\pi 8\widetilde{C}_{3})^{d/2}\frac{1}{(\widetilde{C}_{3}+8t)^{d/2}}\exp{\left(-\frac{|x|^{2}}{\widetilde{C}_{3}+8t}\right)}$ (75)
-since
-$\int_{\mathbb{R}^{d}}e^{-a|z|^{2}}dz=\left(\frac{\pi}{a}\right)^{d/2}.$
-It follows that taking
-$\widetilde{C}_{3}<C_{3}\doteq\min\left(\frac{1}{2\pi};\frac{1}{(8\pi)^{d/2}C}\right),$ (76)
-for $t\in(0,T]$ such that
-$t<T_{1}\doteq\min\left(\frac{\widetilde{C}_{3}}{8};\frac{8(1-2\pi\widetilde{C}_{3})}{2\pi}\right)$ (77)
-we end up with
-$\bar{\rho}(x,t)\leq 2\exp{(-2\pi|x|^{2})}.$ (78)
-3.1.2 Solving (1) for $\sigma\neq 0$
-We assume $\tilde{\rho}$ solves the deterministic PDE
-$\partial_{t}\tilde{\rho}-\Delta\tilde{\rho}+\nabla\cdot(\tilde{\rho}_{t}K\ast\tilde{\rho}_{t})=0$ (79)
-Now we define $\rho(t,x)\doteq\tilde{\rho}(t,x-X_{t})$, where
-$dX_{t}=\sigma_{t}dB_{t},$ (80)
-Hence, by applying the Itô’s formula, since that $\sigma$ does not depends on spatial variable, we find that $\rho$ solves (1), in the sense of Theorem 1.
+When $\sigma=0$, the equation (1) reads
+$\partial_{t}\tilde{\rho}-\Delta\tilde{\rho}+\nabla\cdot(\tilde{\rho}_{t}K\ast\tilde{\rho}_{t})=0,$ (106)
+and a solution to (106), in the sense of Theorem 3, was constructed in sections 3 and 4 of [49] and [50].
+Now, taking into account (19), we define the following process: $\rho(t,x)\doteq\tilde{\rho}(t,x-X_{t})$.
+Hence, by applying the Itô’s formula, since that $\sigma$ does not depends on spatial variable, we find that $\rho$ solves (1), in the sense of Theorem 3, for $\sigma\neq 0$.
 4.2 Inequalities
 In this section, we present several inequalities used throughout the text.
-The first one is the classical Csiszár-Kullback-Pinsker inequality, which is a fundamental result in information theory.
+The first one is the classical Csiszár-Kullback-Pinsker inequality, (see [21]) which is a fundamental result in information theory.
 It quantifies how control of the relative entropy implies control of the total variation distance.
@@
 $\int_{\mathbb{R}^{d}}f\Phi\,dx\leq\eta\mathcal{H}\left(f|g\right)+\eta\ln{\left(\int_{\mathbb{R}^{d}}g\exp{\left(\Phi/\eta\right)}\,dx\right)}.$ (108)
+The next result is a nonlinear Gronwall’s inequality. A proof is provided in Theorem 21 of [17].
+Lemma 6.
+Let $u$ be a nonnegative function satisfying
+$u(t)\leq c(t)+\int_{0}^{t}\big(a(s)\,u(s)+b(s)\,u^{\alpha}(s)\big)\,ds,\qquad c(t)>0,\ \alpha\geq 0,$
+where $a(t)$, $b(t)$ and $c(t)$ are integrable and nonnegative functions on $[0,T]$.
+For $\alpha=1$,
+$u(t)\leq\left(\sup_{t\in[0,T]}c(t)\right)\exp\!\left(\int_{0}^{t}\big(a(s)+b(s)\big)\,ds\right).$ (109)
+For $\alpha>1$ and $c$ a constant function, with the following additional hypothesis, for some $h>0$,
+$c<\left\{\exp\!\left[(1-\alpha)\int_{0}^{h}a(s)\,ds\right]\right\}^{\frac{1}{\alpha-1}}\left\{(\alpha-1)\int_{0}^{h}b(s)\,ds\right\}^{\frac{1}{1-\alpha}},$ (110)
+we get, for $t\in[0,h]$,
+$u(t)\leq c\Biggl\{$ $\exp\left[(1-\alpha)\int_{0}^{t}a(s)\,ds\right]$ (111)
+$-c^{-1}(\alpha-1)\int_{0}^{t}b(s)\exp\left[(1-\alpha)\int_{s}^{t}a(r)\,dr\right]ds\Biggr\}^{\frac{1}{\alpha-1}}.$
 Acknowledgements

### v4

@@
-Quantitative entropy estimates for 2D stochastic vortex model on the whole space under moderate interactions
+Quantitative propagation of chaos for 2D stochastic vortex model on the whole space under moderate interactions
 Abstract
-We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in terms of relative entropy.
-The main novelties lie in the application of the Donsker–Varadhan inequality in the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with the probabilistic data setting, in order to derive estimates for the quadratic variation terms.
-As an application of our entropy bounds, we derive a new energy estimate for the distance between the regularized empirical measure and the solution to the aforementioned model by combining the control of the Fisher information of the particle system, the Ladyzhenskaya inequality, and a nonlinear Grönwall lemma.
-Moreover, we prove the existence of a suitable solution to the limiting process.
+We derive the stochastic 2D vortex model on the whole Euclidean space from moderately interacting particle systems driven by individual and environmental noises, obtaining quantitative estimates in the sense of the entropy and energy functionals.
+The main novelties lie in combining the control of the Fisher information of the particle system with the Ladyzhenskaya and Donsker–Varadhan inequalities, as well as localization techniques within the probabilistic data setting, to address the nonlinearity and quadratic variation arising from Itô’s formula.
+Moreover, we construct a suitable solution for the limiting process.
 MSC2010 subject classification: 35R60, 35Q30, 82C22
 1 Introduction
-In this work, we derive quantitative entropy estimates for the 2D stochastic vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
-$\mathrm{d}\rho_{t}=\Delta\rho_{t}\,dt+\frac{1}{2}D^{2}\rho_{t}(\sigma\sigma^{\top})_{t}\,dt-\nabla\cdot(\rho_{t}(K\ast\rho_{t}))\,dt-\nabla\rho_{t}\cdot\sigma_{t}\,dB_{t},$ (1)
+In this work, we establish a quantitative derivation of the stochastic 2D vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
+$\mathrm{d}\rho_{t}=\Delta\rho_{t}\,dt+\frac{1}{2}\nabla^{2}\rho_{t}(\sigma\sigma^{\top})_{t}\,dt-\nabla\cdot(\rho_{t}(K\ast\rho_{t}))\,dt-\nabla\rho_{t}\cdot\sigma_{t}\,dB_{t},$ (1)
 from a stochastic moderately interacting particle system given by
@@
 where $W_{t}^{i,N}$ and $B_{t}$ are independent standard $\mathbb{R}^{d}$-valued Brownian motions, defined on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$, $K$ is a singular kernel, $V^{N}$ is a suitable scaling and $\sigma$ is a diffusion matrix.
-To achieve these bounds, we study the asymptotic behavior, as $N\to\infty$ of the empirical measure defined by
-$S_{t}^{N}\doteq\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},\qquad t\geqslant 0,$ (3)
-where $\delta_{a}$ is the Dirac delta measure concentrated at $a\in\mathbb{R}^{d}$, by introducing the mollified empirical measure
+To achieve it, we study the asymptotic behavior, as $N\to\infty$ of the empirical measure defined by
+$S_{t}^{N}\doteq\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},\qquad t\geqslant 0$
+by introducing the mollified empirical measure
 $\rho_{t}^{N}\doteq V^{N}\ast S_{t}^{N}=\int_{\mathbb{R}^{d}}V^{N}(\cdot-y)\,S_{t}^{N}(dy),$
-which allows us to apply the relative entropy functional directly to the densities $\rho^{N}$ and $\rho$.
-To the best of our knowledge, this is the first time that pathwise quantitative bounds, in the sense of the relative entropy functional, have been derived for the 2D stochastic vortex model in the whole space via moderately interacting particles.
+which allows us to apply the entropy and energy functionals directly to the densities $\rho^{N}$ and $\rho$.
+Particle systems driven by environmental noises appear in several contexts, such as in statistical mechanics, population dynamics, as well as in numerical analysis, see, e.g., [4], [5], [23], [18], and [27], and quantitative bounds characterize a reduction in the complexity of the system (2), for large values of $N$, see [22].
+To the best of our knowledge, this is the first time that, propagation of chaos, supplemented with pathwise quantitative bounds, in the sense of the relative entropy functional, have been derived for the 2D stochastic vortex model in the whole Euclidean space via moderately interacting particles.
 This extends previously known results for bounded kernels on $\mathbb{R}^{d}$ [29] and for sub-Coulombian kernels under periodic boundary conditions [36].
-As a consequence of our entropy bounds, we obtain energy estimates for the difference between the regularized empirical measure and the solution of (1), extending previous results by [27] and [37].
-Our main novelty lies in the key technical tools we develop.
-For the first time in the context of moderately interacting particles, we apply the Donsker–Varadhan inequality, together with the dissipation provided by the Fisher information, to handle the nonlinearity.
-In addition, we employ localization techniques combined with our probabilistic data setting to derive estimates for the quadratic variation terms in the whole space, since the estimates established in [40] under periodic boundary conditions do not apply in the present framework.
-Finally, new energy estimates quantifying the distance between the regularized empirical measure and the solution to (1) are derived by combining the control of the Fisher information of the particle system with the Ladyzhenskaya inequality and a nonlinear Grönwall inequality.
-Moreover, we prove the existence of a suitable solution to the equation (1).
+In addition, we obtain energy estimates for the difference between the regularized empirical measure and the solution of (1), extending previous results by [25] and [33].
+Our main novelty lies in the key technical tools we combine.
+Using Itô formula, we compute the evolution equations for the entropy and energy functionals associated with the regularized empirical measure and the solution of the aforementioned model.
+By combining the control of the Fisher information of the particle system with the Ladyzhenskaya and Donsker–Varadhan inequalities, along with localization techniques within the probabilistic data setting, to deal with the nonlinearity and quadratic variation, we are able to close the argument with a nonlinear Grönwall inequality.
+Moreover, we prove the existence of a suitable solution to equation (1).
 Related works
 Quantitative estimates for singular interactions have been substantially developed in recent years.
-The relative entropy method for establishing quantitative propagation of chaos in McKean–Vlasov systems with singular interaction kernels was first introduced in [24].
-It was developed in the context of general first-order systems with interaction kernels in $W^{-1,\infty}$, a framework that includes the 2D viscous vortex model.
-We also refer to [43], where quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels are established via the modulated energy method.
+The relative entropy method for establishing quantitative propagation of chaos with singular interaction kernels in $W^{-1,\infty}$ was first introduced in [22], providing a framework that includes the 2D viscous vortex model.
+Alternatively, we refer to [39], which establishes quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels via the modulated energy method.
 In contrast to the relative entropy approach, which operates at the level of the joint law of the particle system, the modulated energy method is formulated directly in terms of the empirical measure associated with the particles.
-Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results.
+Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results, under singular interactions.
 In [2] and [3], the strategies given in [22] and [39], were adapted to obtain a quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes.
@@
 Regarding sub-Coulombian kernels, [6] obtained the derivation of the mean-field approximation for Landau-like equations, while [37] obtained global-in-time mean-field convergence for gradient (conservative) diffusive flows.
-By extending the BBGKY hierarchy introduced in [31] to the case of singular kernels, and adapting the strategy developed in [24], [48] obtained the optimal convergence rate for the relative entropy between the k-particle marginals and the corresponding tensorized law.
+Moreover, quantitative entropy estimates for systems with individual and common noise, have been obtained, as shown in [40] for the 2D vortex model, on torus, [9] for the Hegselmann-Krause model.
 More recently, quantitative propagation of chaos for the 2D vortex model on the whole space was obtained in [44, 45], while the corresponding result for the 2D log-gas was established in [7].
 In these works, the authors exploit Li–Yau-type estimates and Hamilton-type heat kernel estimates to derive suitable bounds for the limiting solution and its derivatives up to second order, which allows them to extend the strategy given in [22], to $\mathbb{R}^{d}$.
-Quantitative estimates in the whole space were derived, in the context of moderately interacting particles  [^2] The moderately interacting particle model was introduced by Oelschläger in [35, 34, 36].
-with singular kernels and without common noises - including the Biot-Savart, the Keller-Segel and the Dirac delta measure kernels - by [37] and [38], based on a semigroup approach introduced in [18], with further advancements in [8], [22], [28] and [39].
+Quantitative estimates in the whole space were derived by [33] and [34] in the context of moderately interacting particles [^2] The moderately interacting particle model was introduced by Oelschläger in [31, 30, 32].
+without common noises and with singular kernels (including the Biot-Savart, Keller-Segel, and Dirac delta measure kernels).
+Their derivation is based on a semigroup approach introduced in [17] and has motivated further advancements, for example, in [8, 15, 20, 26, 35].
+Under periodic boundary conditions, [25] used the Krilov $L^{p}$ theory for SPDE’s, to extend the results given in [33], in the presence of environmental noises, as the semigroup approach is not available in this setting.
 It is important to note that these works do not rely on the relative entropy method, but instead establish quantitative estimates in strong functional topologies.
@@
 However, it was in [36] that equation (1) was derived on the torus for singular kernels and common noise in the moderately interacting regime, by applying the relative entropy directly to the mollified empirical measure.
-Earlier, also in the same domain, the results established in [37] were extended to the setting with environmental noise in [27].
-It is worth highlighting, that quantitative entropy estimates for systems with individual and common noise, have been obtained, as shown in [44] for incompressible the Navier-Stokes equations, [9] for the Hegselmann-Krause model, and [33] for mean-field systems with bounded kernels.
-For more results, see [19], [15], [16], and [29].
-Particle systems driven by environmental noises, such as given by (2), appear in several contexts, such as problems in statistical mechanics, population dynamics and biology, as well as in the analysis of numerical Monte–Carlo schemes and related applications; see, e.g., [4], [5], [25], [20], and [30].
-In contrast to the aforementioned works, our application of the relative entropy method provides pathwise estimates, deriving bounds directly at the trajectory level of the particle system.
-In those works, the estimates are obtained at the level of the joint law.
+In this work, we extend the quantitative entropy estimates given in [36] under periodic boundary conditions and in [29] for bounded kernels to the whole space for the stochastic 2D vortex model.
+Additionaly, we obtain energy estimates for that model, extending the results in [25] and [33].
+In contrast to the aforementioned works, which establish quantitative entropy estimates at the level of the joint law, our application of the relative entropy method provides pathwise estimates, yielding bounds directly at the trajectory level of the particle system.
 Outline of the article
 In order to present our results, this work is organized as follows.
-In the next two subsections, 1.1 and 1.2, respectively, we introduce the notations and assumptions used throughout the text, setting the stage for the subsequent discussions.
+In the next three subsections, 1.1, 1.2, and 1.3, we introduce the notations and assumptions used throughout the text, setting the stage for the subsequent discussions, as well as the notion of solution to (1).
 In the subsection 1.4, we present our main results.
 The sections 2 and 3 are dedicated to the proofs of our main results.
-To ease the presentation, the Appendix, provides the existence of a suitable solution for the limiting equation, as well as some inequalities used throughout the proofs.
+Finally, the Appendix provides the existence of a suitable solution for the limiting equation, as well as some inequalities used throughout the proofs.
 1.1 Notations
@@
   1. $(\mathbf{A}^{V})$
-    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)=N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V\in\mathcal{P}\left(\mathbb{R}^{d}\right)$ satisfying
+    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)\doteq N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V\in\mathcal{P}\left(\mathbb{R}^{d}\right)$ satisfying
     $\left|\nabla V\right|\leq C_{d}V,$ (4)
     for $\alpha>d/2$
-    $V(y)\leq C_{d}(1+|y|^{2})^{-\alpha},$ (6)
-    and for $\gamma\in(0,1)$ and $y\neq 0$
-    $V(y)\leq C_{d}|y|^{-3d-\gamma}.$ (7)
+    $V(y)\leq C_{d,\alpha}(1+|y|^{2})^{-\alpha},$ (5)
+    and for $\gamma,\beta\in(0,1)$ and $y\neq 0$
+    $V(y)\leq C_{d,\gamma,\beta}|y|^{-3d/\beta-\gamma}.$ (6)
   2. $(\mathbf{A}^{K})$
-    It holds for $\gamma\in(1/2,1)$ that
-    $\|K\ast f\|_{\gamma}\leq C\left(\|f\|_{1}+\|f\|_{4}\right)$ (8)
-    for all $f\in H^{1}$.
+    It holds for $\gamma\in(0,1/2)$ that
+    $\|K\ast f\|_{\gamma}\leq C_{\gamma}\left(\|f\|_{1}+\|f\|_{4}\right)$ (7)
+    for all $f\in W^{1,2}$.
   3. $(\mathbf{A}^{\nabla\cdot K})$
@@
   5. $(\mathbf{A}^{\rho_{0}})$
-    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,\infty}\cap\mathcal{P}\left(\mathbb{R}^{d}\right)$ and
+    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,1}\cap W^{2,\infty}\cap\mathcal{P}\left(\mathbb{R}^{d}\right)$, and
     $\left|\nabla\ln{\rho_{0}}(x)\right|^{2}\leq\widetilde{C}_{1}(1+|x|^{2}),$ (9)
@@
 $K_{0}(x_{1},x_{2})=-\left(\frac{1}{2\pi}\arctan{\left(\frac{x_{1}}{x_{2}}\right)}\right)I.$ (13)
-It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$. Furthermore, $K$ satisfy $(\mathbf{A}^{K})$, see [39].
+It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$. Furthermore, $K$ satisfies $(\mathbf{A}^{K})$, see [35], page 8.
 Remark 3.
@@
 It allows the relative entropy approach of [22] to be extended from the periodic framework to the whole space.
-1.3 Statement of the main results
-In this subsection, we establish our main results. We begin by formulating a suitable notion of solution, needed for deriving quantitative entropy estimates.
+It motivates the following definition.
+1.3 Solution of limiting process
+We now formulate a suitable notion of solution of (1) to derive our quantitative propagation of chaos.
 Definition 1.1.
 Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$.
-A pathwise solution for (1) with initial condition $\rho_{0}$, is a measurable function $\rho:[0,T_{1}]\times\mathbb{R}^{d}\times\Omega\to\mathbb{R}$, for some $T_{1}>0$, such that:
+A pathwise solution for (1) with initial condition $\rho_{0}$ is a measurable function $\rho:[0,T]\times\mathbb{R}^{d}\times\Omega\to\mathbb{R}$, for some $T>0$, such that:
   • it holds $\mathbb{P}$-a.s.
-    $\rho\in L^{\infty}\left([0,T_{1}];W^{2,1}\cap W^{2,\infty}\cap\mathcal{P}(\mathbb{R}^{d})\right)$ (15)
+    $\rho\in L^{\infty}\left([0,T];W^{2,1}\cap W^{2,\infty}\cap\mathcal{P}(\mathbb{R}^{d})\right)$ (14)
   • the solution $\rho$ verifies, $\mathbb{P}$-a.s., that
@@
     $dX_{t}=\sigma_{t}dB_{t},$ (18)
-    for all $t\in[0,T_{1}]$ and constants $C_{1},C_{2}$ and $C_{3}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
-We are now able to extend what was known before for bounded kernels in the whole space as in [33] and in [40], for sub-Coulombian kernels, under periodic boundary conditions.
-Our main result can be stated as follows.
+    for all $t\in[0,T]$ and constants $C_{1},C_{2}$ and $C_{3}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
+1.4 Statement of the main results
+We are now able to extend what was known before for bounded kernels in the whole space as in [29] and in [36], for sub-Coulombian kernels, under periodic boundary conditions, by deriving entropy estimates.
+Also we extend previous results by [25] and [33], by establishing energy estimates.
+Our main result is stated as follows.
+Theorem 1.
+Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{K})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$, $d=2$ and let $\rho$ be a solution of (1), given by Definition 1.1.
+In addition, let the dynamics of the particle system be given by (2) and
+$\lim_{N\to\infty}N^{\theta}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.,$ (19)
+$\lim_{N\to\infty}N^{\theta}\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}=0,\,\,\,\mathbb{P}-a.s.$ (20)
+and
+$\theta\doteq\min{\left(\left(1-\beta\left(2+2\alpha\right)\right);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right);\frac{\beta\gamma^{2}}{2}\right)}-\delta$
+with $\gamma\in(0,1/2)$ and $\delta>0$, such that $\theta>0$ and $\beta\in\left(0,\frac{1}{4}\right)\cap\left(0,\frac{1}{\big[2+2\alpha\big]}\right)$.
+Then, there exists a time $T>0$, such that
+$\lim_{N\to\infty}N^{\theta}\left(\sup_{t\in[0,T]}\left(\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}\right)+\int_{0}^{T}\left\|\nabla(\rho_{s}^{N}-\rho_{s})\right\|_{2}^{2}\,ds\right)$ $=0,$ (21)
+almost surely.
+As a corollary of Theorem 1, we establish quantitative propagation of chaos for the marginals of the empirical measure of the particle system.
+In the sense of [1, Section 8.3], it corresponds to get convergence of the empirical measure in the sense to the Kantorovich–Rubinstein metric, given in (3).
+Corollary 1.
+Under the assumptions of Theorem 1, we have
+$\lim_{N\to\infty}N^{\theta}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $=0,\,\,\,\mathbb{P}-a.s.$ (22)
+In dimensions greater than or equal to two, we establish the boundedness of the relative entropy functional:
 Theorem 2.
@@
 and
-$\theta=\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
+$\theta=\min\left(1-\beta\left(1+\frac{2}{d}+2\alpha\right);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
 with $\delta>0$, such that $\theta>0$, $d\geq 1$ and $\beta\in\left(0,\frac{1}{2\big[1+\frac{1}{d}\big]}\right)\cap\left(0,\frac{1}{\big[1+\frac{2}{d}+2\alpha\big]}\right)$.
-Then, there exist a time $T\in(0,T_{1})$, an almost surely finite random variable $A_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
-$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (21)
+Then, there exist a time $T>0$, an almost surely finite random variable $A_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
+$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (24)
 In particular, we get
 $\limsup_{N\to\infty}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim T,\,\,\,\,\,\,\mathbb{P}-a.s.$ (25)
-The following result establishes energy estimates as an application of our entropy bounds, extending previous results by [27] and [37].
-Precisely, we utilize the entropy dissipation to control the Fisher information of the particle system, which in turn allows us to derive, a new energy estimate for the difference between the regularized empirical measure and the solution to (1), using the Ladyzhenskaya inequality and a nonlinear Gronwall lemma.
-Theorem 2.
-Assume $(\mathbf{A}^{V})$, $(\mathbf{A}^{K})$, $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, $(\mathbf{A}^{\rho_{0}})$, $(\mathbf{A}^{X_{0}})$, $d=2$ and let $\rho$ be a solution of (1), given by Definition 1.1.
-In addition, let the dynamics of the particle system be given by (2) and
-$\lim_{N\to\infty}N^{\widetilde{\theta}}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.,$ (23)
-$\lim_{N\to\infty}N^{\widetilde{\theta}}\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}=0,\,\,\,\mathbb{P}-a.s.$ (24)
-and
-$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
-where
-$\theta_{1}\doteq\left(1-\beta\left(2+2\alpha\right)\right);\,\,\,\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right);\,\,\,\widetilde{\theta}_{1}\doteq\left(1-2\beta(1+\frac{1}{2})+\beta\right)$
-$\widetilde{\theta}_{2}\doteq\beta(2-3\gamma);\,\,\,\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{2};\,\,\,\widetilde{\theta}_{4}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right)$
-with $\gamma\in(1/2,2/3)$ and $\delta>0$, such that $\widetilde{\theta}>0$ and $\beta\in\left(0,\frac{1}{2+2\alpha}\right)$.
-Then, there exist a time $T\in(0,T_{1})$, a pair of almost surely finite random variables $A_{0}$, $\widetilde{A}_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
-$\sup_{t\in[0,T]}\|\rho_{t}^{N}-\rho_{t}\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
-[…]
-In particular, we get
-$\limsup_{N\to\infty}\left(\sup_{t\in[0,T]}\|\rho_{t}^{N}-\rho_{t}\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds\right)$ $\lesssim T,\,\,\,\mathbb{P}-a.s.$ (26)
-As a corollary of Theorem 1, we establish quantitative bounds for the marginals of the empirical measure of the particle system.
-In the sense of [1, Section 8.3], it corresponds to get bounds of the empirical measure in the sense to the Kantorovich–Rubinstein metric, given in (4).
-Corollary 1.
-Under the assumptions of Theorem 1, for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$, such that for $N\geq N_{0}$ we have
-$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (27)
-In particular, we get
-$\limsup_{N\to\infty}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim T,\,\,\,\mathbb{P}-a.s.$ (28)
-Remark 4.
-We observe that, compared with previous works on moderate particle systems without common noise, we lose the order of convergence.
-This is because the estimation of the martingale term in our work is is of order $\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)$, while in [37] is the order $\frac{1}{2}-\frac{\beta}{2}$, see Theorem 1.3 in [37].
-When environmental noise is taken into account, the overall framework remains comparable to that of [27].
-For the same reason, the range of the parameter $\beta$ is smaller in our work.
-However, in the Theorem 1, we obtain an improved convergence rate compared to that established in [40].
 The next result, whose proof is postponed to the Appendix, establishes that the limiting equation is well posed in the appropriate functional framework, given by Definition 1.1.
@@
 Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. Then, there exists a solution of (1) in the sense of Definition 1.1.
-2 Proofs of Theorem 1
+2 Proof of Theorem 1
 In this section we present the proofs of Theorem 1.
 In Subsection 2.1, by applying the Itô’s formula, such as in [36], we derive an evolution equation for the relative entropy functional of the regularized empirical measure with respect to the solution of (1).
-In Subsection 2.2, we deal with the quadratic variation terms arising from the application of Itô’s formula.
-In Subsection 2.3, we address the nonlinear contributions appearing in the evolution equation for the relative entropy functional obtained in the first step.
-Finally, in Subsections 2.4 and 2.5, combining the previous estimates with Gronwall’s lemma and suitable localization techniques, we close the argument and complete the proof.
+In Subsection 2.1.1, we deal with the quadratic variation terms arising from the application of Itô’s formula, while in Subsection 2.1.2, we address the nonlinear contributions appearing in the evolution equation for the relative entropy functional obtained in the first step.
+In Subsection 2.2, we derive an evolution equation for the energy, relating our densities, being the nonlinear term, arising from this computation, estimated in Subsection 2.2.1, as in Subsection 2.3, we combine our entropy and energy estimates.
+Finally, in Subsection 2.4, by a nonlinear Gronwall Lemma and suitable localization techniques, we close the argument and complete the proof.
 2.1 Time evolution of the relative entropy
@@
 $-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds,$ (27)
-$II_{t}+IV_{t}$ $=\frac{1}{2N^{2}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{1}{\rho_{s}^{N}}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx,$ (31)
-$III_{t}$ $=-\frac{1}{2}\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds,$ (32)
+$II_{t}+IV_{t}$ $=\frac{1}{N^{2}}\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{1}{\rho_{s}^{N}}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,dx\,ds,$ (28)
+$III_{t}$ $=-\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds,$ (29)
 and
-$M_{t}^{N}$ $=\int_{\mathbb{R}^{d}}\int_{0}^{t}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dB_{s}\,dx$
-$+\int_{\mathbb{R}^{d}}\int_{0}^{t}\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dB_{s}\,dx$
-$+\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\nabla V^{N}(x-X_{s}^{i})\,dW^{i}_{s}\,dx$
-$-\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\nabla V^{N}(x-X_{s}^{i})\,dW^{i}_{s}\,dx$
-$+\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{\rho_{s}^{N}}{\rho_{s}}\sigma^{\top}_{s}\nabla\rho_{s}\,dB_{s}\,dx.$ (33)
-2.2 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
+$M_{t}^{N}$ $=\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dx\,dB_{s}$
+$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dx\,dB_{s}$
+$+\frac{1}{N}\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\nabla V^{N}(x-X_{s}^{i})\,dx\,dW^{i,N}_{s}$
+$-\frac{1}{N}\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla V^{N}(x-X_{s}^{i})\,dx\,dW^{i,N}_{s}$
+$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{N}}{\rho_{s}}\sigma^{\top}_{s}\nabla\rho_{s}\,dx\,dB_{s}.$ (30)
+2.1.1 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
 We now derive estimates for the quadratic variation terms appearing in (26).
@@
 First, taking into account Assumption $(\mathbf{A}^{V})$ in (28), we have
-$II_{t}+IV_{t}$ $\leq\frac{N^{2\beta/d}}{2N^{2}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{1}{\rho_{s}^{N}}|V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx.$ (34)
+[…]
 We introduce the following stopping time:
 $\tau^{N}=\inf\left\{t\geq 0;\exists i\in\{1,...,N\},|X_{t}^{i,N}|\geq N^{\beta}\right\}\wedge T.$ (32)
-By decay of $\rho_{0}$ in (12), along with Borel-Cantelli Lemma, almost surely, there exists $N_{0}=N_{0}(\omega)$ such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$.
+By decay of $\rho_{0}$ in (11), along with Borel-Cantelli Lemma, there exists $\Lambda_{0}\in\mathcal{F}$, $\mathbb{P}(\Lambda_{0})=1$, with the following property: for all $\omega\in\Lambda_{0}$, there exists $N_{0}=N_{0}(\omega)$ such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$.
 Indeed, setting
@@
 where in the second inequality we are using (11) to get
-$\int_{|x|\geq N^{\beta}}\rho_{0}(x)dx$ $\leq\widetilde{C}_{3}\int_{|x|\geq N^{\beta}}e^{-\widetilde{C}_{3}^{-1}|x|^{2}}dx$
+[…]
 $=C_{d}\int_{r=N^{\beta}}^{\infty}e^{-\widetilde{C}_{3}^{-1}r^{2}}r^{d-1}dr\leq C_{d}(N^{\beta})^{d-2}e^{-\widetilde{C}_{3}^{-1}(N^{\beta})^{2}}.$ (35)
@@
 $\mathbb{P}\left(\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}\Lambda^{N}\right)=0.$ (36)
-With this in view for all $\omega\in\Lambda_{0}$, with
-$\Lambda_{0}\doteq\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}$ (40)
-there exists $N_{0}=N_{0}(\omega)$, such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$, since $t\mapsto X_{t}^{i,N}(\omega)$ is continuous, almost surely.
+With this in view, since $t\mapsto X_{t}^{i,N}$ is continuous, for all $i\in\{1,...,N\}$, almost surely, by setting
+$\Lambda_{X_{0}}\doteq\bigcap_{N\in\mathbb{N}}\bigcap_{i=1}^{N}\left\{t\mapsto X_{t}^{i,N}\,\,\,\text{is continuous}\right\}$
+and
+$\Lambda_{0}\doteq\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}\right)\cap\Lambda_{X_{0}}$ (37)
+we have that $\mathbb{P}(\Lambda_{0})=1$, and for all $\omega\in\Lambda_{0}$, there exists $N_{0}=N_{0}(\omega)$, such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$.
 Let $s\in(0,\tau^{N})$. We observe that,
@@
 $+\int_{|x|>2|X_{s}^{i,N}|}\frac{|V^{N}(x-X_{s}^{i,N})|^{2}}{\rho^{N}_{s}(x)}\,dx.$ (38)
-Now, if $|x|\leq 2|X_{s}^{i,N}|$, then $|x|\leq 2N^{\beta}$ for $s\in(0,\tau^{N})$, and since $V^{N}\leq N^{\beta}$, we derive for $\alpha>d/2$
+Now, if $|x|\leq 2|X_{s}^{i,N}|$, then $|x|<2N^{\beta}$ for $s\in(0,\tau^{N})$, and since $V^{N}\leq N^{\beta}$, we derive for $\alpha>d/2$
 $V^{N}(x-X_{s}^{i,N})\leq N^{\beta}\left(1+|x|^{2}\right)^{\alpha-\alpha}\leq C_{\alpha}N^{2\alpha\beta+\beta}\left(1+|x|^{2}\right)^{-\alpha}.$ (39)
@@
 with
-$\phi(x)\doteq\left(\left(1+|x|^{2}\right)^{-\alpha}+\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\right)$ (45)
+$\phi(x)\doteq\left(\left(1+|x|^{2}\right)^{-\alpha}+\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\right).$ (42)
 Hence by (41) in (31) we obtain
 […]
-$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
-$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
-$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
-$\overset{\alpha>d/2}{\leq}C_{\alpha,T}N^{-\theta_{1}},$ (46)
+$\overset{}{\leq}C_{\alpha,d}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
+$\overset{}{\leq}C_{\alpha,d}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
+$\overset{}{\leq}C_{\alpha,d}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
+$\overset{\alpha>d/2}{\leq}C_{\alpha,d,T}N^{-\theta_{1}},$ (43)
 for $\theta_{1}\doteq\left(1-\beta\left(1+2/d+2\alpha\right)\right)$.
-2.3 Estimates for $I_{t}$: the nonlinear terms
+2.1.2 Estimates for $I_{t}$: the nonlinearity
 We now address the nonlinear terms in (26).
-The main novelty is to use Donsker-Varadhan inequality (108), combined with the notion of solution of limiting process given by Definition 1.1, in the context of moderate interactions, along with the dissipation that comes from of Fisher information and Assumption $(\mathbf{A}^{\nabla\cdot K})$, to tackle with the singularity of kernel $K$.
+The main novelty is to use the stopping time obtained in previous Subsection, along with the dissipation that comes from of Fisher information and Ladyzenkaya inequality, to tackle with the singularity of kernel $K$.
 First, we set
 $I_{t}$ $\overset{}{=}\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho_{s}^{N}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)[K\ast\rho_{s}-K\ast\rho_{s}^{N}]\,dx\,ds$
-$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\underbrace{\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle}_{\doteq I^{N}}\,dx\,ds$
-$\doteq I^{1}_{t}-I^{2}_{t},$ (47)
-where
-$I^{1}_{t}$ $=-\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\rho_{s}^{N}\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
-$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\nabla\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
-$\doteq-I^{1,1}_{t}-I^{1,2}_{t}.$ (48)
-Now by $\epsilon$-Young inequality and convolution inequality, we deduce
-$I_{t}^{1,2}$ $=\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\nabla\rho_{s}\left(\frac{\sqrt{\rho_{s}^{N}}}{\rho_{s}}\right)\left(\frac{\rho_{s}}{\sqrt{\rho_{s}^{N}}}\right)\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
-$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\left|K_{0}\ast(\rho_{s}-\rho_{s}^{N})\right|^{2}\,dx\,ds$
-$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\|K_{0}\|_{\infty}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\,dx\,ds.$ (49)
-We observe that, for $\eta>2C_{1}(8T+C_{3})+1$
-$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty$ (50)
-which implies by Donsker-Varadhan inequality (108),
+$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
+$\doteq I_{t}^{1}+I_{t}^{2}$ (44)
+Thus, by Young inequality together with (7), we note that
+$I_{t}^{1}$ $\leq\epsilon\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon}\int_{0}^{t}\left(\|\rho_{s}^{N}-\rho_{s}\|_{1}^{2}+\|\rho_{s}^{N}-\rho_{s}\|_{4}^{2}\right)\,ds$
+$\leq\epsilon\int_{0}^{t}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+\|\nabla(\rho^{N}_{s}-\rho_{s})\|_{2}^{2}\right)\,ds$
+$\hskip 80.29999pt+C_{\epsilon,d}\int_{0}^{t}\left(\|\rho_{s}^{N}-\rho_{s}\|_{1}^{2}+\|\rho_{s}^{N}-\rho_{s}\|_{2}^{2}\right)\,ds,$ (45)
+where, in the second step, we use the Ladyzhenskaya and Young inequalities.
+Now, we will focus on the term $I_{t}^{2}$ in (44). First, we observe that by subtracting and adding $K\ast\rho_{s}(\cdot)$ and $K\ast\rho_{s}(x)$ along with (7)
+$\left|K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right|$ $\lesssim\left(\|\rho^{N}_{s}-\rho_{s}\|_{1}+\|\rho^{N}_{s}-\rho_{s}\|_{4}\right)$
+$\hskip 80.29999pt+\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|.$ (46)
+Additionally, by (7) and (17), we obtain
+$\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|\right\rangle$
 […]
 […]
-Hence by Csiszár-Kullback-Pinsker inequality (107) and (49), we derive
-$I_{t}^{1,2}$ $\leq\epsilon\int_{0}^{t}\mathcal{I}(\rho^{N}_{s}|\rho_{s})\,ds+C_{\epsilon,\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds.$ (52)
-Regarding term $I_{t}^{1,1}$ in (48), by convolution inequality, we find
-$I_{t}^{1,1}$ $=\int_{0}^{t}\|K_{0}\|_{\infty}\|\rho^{N}_{s}-\rho_{s}\|_{1}\int_{\mathbb{R}^{d}}\left|\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right|\rho^{N}_{s}\,dx\,ds.$ (53)
-In addition, if $C_{1,2}\doteq\max{(C_{1},C_{2})}$ since
+We define
+$R\doteq\left\{(x,\cdot);|x-\cdot|<N^{-\beta\gamma/d}\right\}$ (48)
+It follows by (47) and (48) that
+$\int_{\mathbb{R}^{d}}\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|\mathds{1}_{R}(x,\cdot)\right\rangle\lesssim N^{-\beta\gamma^{2}/d}.$ (49)
+On the other hand, by decay of $V$, given by (6), we have
+$V^{N}(x-\cdot)|x-\cdot|^{\gamma}$ $=N^{\beta}V(N^{\beta/d}(x-\cdot))|x-\cdot|^{\gamma}$
 […]
-for $\eta>2C_{1,2}(8T+C_{3})+1$, we obtain
-$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty.$ (55)
-It follows by Donsker-Varadhan inequality (108),
+$\lesssim N^{-3+\beta}\left|(x-\cdot)\right|^{-3d/\beta}.$ (50)
+Hence, we derive
+$\int_{\mathbb{R}^{d}}\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\,dx$
 […]
 […]
-Therefore, by (56) in (53), we get
-$I_{t}^{1,1}$ $\overset{}{\leq}C_{\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (57)
+$\hskip-170.6375pt+N^{-3+\beta}\int_{N^{\beta}<|x|\leq 2N^{\beta}}\left|\left\langle S_{s}^{N},\left|(x-\cdot)\right|^{-3d/\beta}\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\right|\,dx$
+$\hskip-170.6375pt+N^{-3+\beta}\int_{|x|>2N^{\beta}}\left|\left\langle S_{s}^{N},\left|(x-\cdot)\right|^{-3d/\beta}\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\right|\,dx$
+$\hskip-90.3375pt\doteq I^{1}+I^{2}+I^{3}.$ (51)
+Now, for $\gamma\in(0,1/2)$, we get
+$I^{1}$ $=N^{-3+\beta}\int_{|x|\leq N^{\beta}}\left|\left\langle S_{s}^{N},\left|(x-\cdot)\right|^{-3d/\beta}\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\right|\,dx\leq N^{-3+\beta}N^{d\beta}N^{3\gamma}$
+$\leq N^{-3+(1+d)\beta+3\gamma}.$ (52)
+Also, for the same reason we arrive at,
+$I^{2}$ $=N^{-3+\beta}\int_{N^{\beta}<|x|\leq 2N^{\beta}}\left|\left\langle S_{s}^{N},\left|(x-\cdot)\right|^{-3d/\beta}\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\right|\,dx\leq C_{d}N^{-3+\beta}N^{d\beta}N^{3\gamma}$
+$\leq C_{d}N^{-3+(1+d)\beta+3\gamma}.$ (53)
+Moreover, if $|x|>2N^{\beta}$, since that for $s\in(0,\tau^{N})$, $|X_{s}^{i,N}|<N^{\beta}$ (see (33)), for all $i\in\{1,...,N\}$, we have, $|x-\cdot|\geq 1/2|x|$, which implies
+$\int_{|x|>2N^{\beta}}|x-\cdot|^{-3d/\beta}\,dx\leq C_{d}\int_{|x|>2N^{\beta}}|x|^{-3d/\beta}\,dx=C_{d}N^{-3d+d\beta}$
+and then,
+$I^{3}$ $=N^{-3+\beta}\int_{\mathbb{R}^{d}}\int_{|x|>2N^{\beta}}|x-y|^{-3d/\beta}\mathds{1}_{R^{c}}(x,y)\,dx\,S_{s}^{N}(dy)$
+$\leq N^{-3+\beta}C_{d}(N^{\beta})^{-3d/\beta+d}\leq C_{d}N^{-3+(1+d)\beta+3\gamma}.$ (54)
+It follows by (51)-(54) that,
+$\int_{\mathbb{R}^{d}}\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\lesssim N^{-3+(1+d)\beta+3\gamma}.$ (55)
+Therefore, by (49) and (55), $\theta_{2}\doteq\beta\gamma^{2}/d$ with $\gamma\in(0,1/2)$, and $s\in(0,\tau^{N})$, we deduce,
+$\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right|\right\rangle\,dx$
 […]
-$I_{t}^{1}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\epsilon\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (58)
-Now, we will focus on the term $I_{t}^{2}$ in (47). First, we recall that by Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$, $K=\nabla\cdot K_{0}$ with $K_{0}\in L^{\infty}$. Hence, we arrive at
-$\left|K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right|$ $=\left|K_{0}\ast\nabla\rho_{s}^{N}(\cdot)-K_{0}\ast\nabla\rho_{s}^{N}(x)\right|$
-$\leq\int_{\mathbb{R}^{d}}\left|K_{0}(\cdot-y)-K_{0}(x-y)\right||\nabla\rho_{s}^{N}(y)|\,dy$
-$\leq\int_{\mathbb{R}^{d}}2\|K_{0}\|_{\infty}|\nabla\rho_{s}^{N}(y)|\,dy$
-$=2\|K_{0}\|_{\infty}\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}(y)|\,dy.$ (59)
-Additionally, $\rho_{s}^{N}\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, $\mathbb{P}$-a.s. and then by Holder’s inequality we obtain
-$\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\,dx$ $=\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\frac{\sqrt{\rho_{s}^{N}}}{\sqrt{\rho_{s}^{N}}}\,dx$
-$\overset{\text{Holder}}{\leq}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}^{d}}\rho_{s}^{N}\,dx\right)^{\frac{1}{2}}$
-$\overset{}{=}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}.$ (60)
-Now from Leibniz’s rule we deduce,
-$\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx$ $=\int_{\mathbb{R}^{d}}\frac{\left|\nabla\left(\rho_{s}\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
-$\leq 2\int_{\mathbb{R}^{d}}\frac{\left|\nabla\rho_{s}\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx+2\int_{\mathbb{R}^{d}}\frac{\left|\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
-$\overset{}{\leq}2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\frac{\left|\nabla\rho_{s}\right|^{2}}{\rho_{s}^{2}}\,dx+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx$
+$+\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(\rho_{s}^{N})^{1/2}\left|\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)]\right\rangle\right|^{1/2}\,dx$
 […]
 […]
-$I^{N}$ $\leq 2\|K_{0}\|_{\infty}\left\langle S_{s}^{N},V^{N}(x-\cdot)\right\rangle\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}$
-$\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}},$
-which implies, by $\epsilon$-Young and Holder inequalities,
-$I^{2}_{t}$ $\doteq\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
-$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
-$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})\right)^{\frac{1}{2}}\,dx\,ds$
-$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\,dx\,ds$
-$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
-$\overset{\text{Young}}{\leq}\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds$
-$+\sqrt{2}/2\int_{0}^{t}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\,dx\,ds$
-$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\epsilon,\eta}$
-$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\epsilon,\eta}.$ (62)
-[…]
-$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\epsilon,\eta}.$ (63)
-2.4 Application of Gronwall’s Lemma
-In this subsection, we conclude our localizated estimates, by applying Gronwall’s inequality.
-Indeed, with the same computations as in [40], for all $\delta\in(0,\theta_{2})$ there exists a random variable $A_{0}$, finite almost surely, such that
-$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{2}+\delta},$ (64)
-with $\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
-[…]
-$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds$
-$\overset{}{+}(3\epsilon+\sqrt{2}/2)\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
-$-\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
-$+C_{\alpha,T}N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+TC_{\eta}$ (65)
-and then for $\eta>>1$ and $\epsilon<<1$, we deduce
-$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+T,$ (66)
-with $\theta_{1}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$ and $\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
-Finally, from Gronwall’s inequality (109) in (66), recalling (36), (39) and (40), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
-$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right)$ (67)
-with
-$\theta\doteq\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta\right).$
-2.5 End of the proof: Removing the stopping time
-We now remove the stopping time $\tau^{N}$ introduced in (67) by exploiting the probabilistic framework provided by Assumptions $\left(\mathbf{A}^{\rho_{0}}\right)$ and $\left(\mathbf{A}^{X_{0}}\right)$.
-It allows us to pass from the localized estimates to uniform bounds with high probability, removing the stopping time.
-First, since $K=\nabla\cdot K_{0}$, $K_{0}\in L^{\infty}$ and $V^{N}=N^{\beta}V(N^{\beta/d}\cdot)$, by (2), we have
-$\sup_{t\in[0,T]}|X_{t}^{i,N}|$ $\leq|X_{0}^{i,N}|+T\|\nabla V\|_{1}\|K_{0}\|_{\infty}N^{\beta}$
-$+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sqrt{2}dW_{s}^{i,N}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|,$ (68)
-which implies for $\widetilde{C}_{K_{0},V}\doteq(1-T\|\nabla V\|_{1}\|K_{0}\|_{\infty})$
-$\{\tau^{N}<T\}$ $\subset\bigcup_{i=1}^{N}\left\{\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
-$=\left\{\exists i;\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
-$\subset\left\{\exists i;|X_{0}^{i,N}|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sqrt{2}dW_{s}^{i,N}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|\geq\widetilde{C}_{K_{0},V}N^{\beta}\right\}$
-$\subset L_{1}^{N}\cup L_{2}^{N}\cup L_{3}^{N},$ (69)
-with
-$L_{1}^{N}\doteq\left\{\exists i;|X_{0}^{i,N}|\geq C_{K_{0},V}N^{\beta}\right\},$ (70)
-$L_{2}^{N}\doteq\left\{\exists i;\sup_{t\in[0,T]}\left|\int_{0}^{t}dW_{s}^{i,N}\right|\geq C_{K_{0},V}N^{\beta}\right\}$ (71)
-and
-$L_{3}^{N}\doteq\left\{\exists i;\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|\geq C_{K_{0},V}N^{\beta}\right\}$ (72)
-with $C_{K_{0},V}\doteq(1-T\|\nabla V\|_{1}\|K_{0}\|_{\infty})/3\sqrt{2}>0$, for $T<(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})$.
-We note that, with the same computations like those led (37), we deduce
-$\mathbb{P}\left(L_{1}^{N}\right)$ $\overset{}{\leq}N(C_{K_{0},V}N^{\beta})^{d-2}e^{-(C_{K_{0},V}N^{\beta})^{2}}$ (73)
-We now take care about the $L_{2}^{N}$. Indeed, we note that $W_{t}^{i,N}=\int_{0}^{t}dW_{s}^{i,N}$, and since
-$|W_{t}^{i,N}|^{2}=\sum_{j=1}^{d}(W_{t}^{i,N,j})^{2},$
-if $|W_{t}^{i,N}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}.$ Therefore,
-$\left\{\sup_{t\leq T}|W_{t}^{i,N}|\geq a\right\}\subset\bigcup_{j=1}^{d}\left\{\sup_{t\leq T}|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}\right\},$ (74)
-we obtain
-[…]
-$\leq\sqrt{\frac{T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}$
-$\times\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2T)\right)},$ (75)
-where in the second inequality, we are using (8.3)’ in Chapter 2 of [26].
-Now we will derive an analogous estimate for $L_{3}^{N}$.
-First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$.
-Since $\sigma$ is bounded, for $j\in\{1,...,d\}$, the function defined by $t\mapsto\langle X^{j}\rangle_{t}$ is continuous and increasing.
-Hence, recalling (19), by Theorem 4.6, in Chapter 3 of [26], for all $j\in\{1,...,d\}$, we get
-$\sup_{t\leq T}|X_{t}^{j}|=\sup_{t\leq T}\left|\widetilde{B}^{j}_{\langle X^{j}\rangle_{t}}\right|\leq\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j}_{t}|,$ (76)
-where $(\widetilde{B}^{j}_{t})_{t\geq 0}$ is the time-change process, which in particular, is a one dimensional standard Brownian motion. In addition, since
-$|X_{t}|^{2}=\sum_{j=1}^{d}|X_{t}^{j}|^{2}$
-if $|X_{t}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|X_{t}^{j}|\geq\frac{a}{\sqrt{d}}.$ Therefore, for all $a>0$, we find
-[…]
-It follows that, by (76) and (77)
-[…]
-$\leq\sqrt{\frac{\|\sigma\|_{\infty}^{2}T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}$
-$\times\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2\|\sigma\|_{\infty}^{2}T)\right)},$ (78)
-where in the second inequality, we are using (8.3)’ in Chapter 2 of [26].
-Hence, by (73), (75) and (78) in (69), we find
-$\sum_{N\in\mathbb{N}}\mathbb{P}\left(\tau^{N}<T\right)<\infty$ (79)
-and then by Borel-Cantelli Lemma, we arrive at
-$\mathbb{P}\left(\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}\{\tau^{N}<T\}\right)=0.$ (80)
-Finally, by (40) and (80), for $T<\min{\left(T_{1};(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})\right)}$ for all $\omega\in\Lambda$, with
-$\Lambda\doteq\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}\right)\cap\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}\{\tau^{N}=T\}\right)\cap\left\{A_{0}<\infty\right\},$ (81)
-there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, we have $\tau^{N}(\omega)=T$ and
-$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (82)
-In particular, by (20) and (64), we conclude with
-$\limsup_{N\to\infty}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim T,$ (83)
-almost surely.
-3 Proof of Theorem 2
-In this section we present the proofs of Theorem 2.
-In Subsection 3.1, by applying the Itô’s formula, such as in [27], we derive an evolution equation for the energy of the difference between regularized empirical measure and the solution of (1).
-In Subsection 3.2, we address the nonlinear terms, combining bounds on the Fisher information of the particle system with the Ladyzhenskaya inequality.
-Finally, in Subsections 3.3, combining the previous estimates with a nonlinear Gronwall’s lemma, we complete the proof.
+$\hskip 120.45pt+C_{\epsilon,\gamma,\beta,d}\left(\|\rho^{N}_{s}-\rho_{s}\|_{1}^{2}+\|\rho^{N}_{s}-\rho_{s}\|_{2}^{2}+N^{-\theta_{2}}\right),$ (56)
+where the last step follows from the Ladyzhenskaya and Young inequalities.
+Finally, by (56), we get
+$I_{t\wedge\tau^{N}}^{2}$ $\leq\int_{0}^{t\wedge\tau^{N}}\epsilon\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+\|\nabla(\rho_{s}^{N}-\rho_{s})\|_{2}^{2}\right)\,ds$
+$\hskip 80.29999pt+C_{\epsilon,\gamma,\beta,d}\int_{0}^{t\wedge\tau^{N}}\left(\|\rho^{N}_{s}-\rho_{s}\|_{1}^{2}+\|\rho^{N}_{s}-\rho_{s}\|_{2}^{2}+N^{-\theta_{2}}\right)\,ds,$ (57)
+and thus, by (44), (45) and (57), we end up with
+$I_{t\wedge\tau^{N}}$ $\leq 2\epsilon\int_{0}^{t\wedge\tau^{N}}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+\|\nabla(\rho_{s}^{N}-\rho_{s})\|_{2}^{2}\right)\,ds$
+$\hskip 80.29999pt+C_{\epsilon,\gamma,\beta,d}\int_{0}^{t\wedge\tau^{N}}\left(\|\rho^{N}_{s}-\rho_{s}\|_{1}^{2}+\|\rho^{N}_{s}-\rho_{s}\|_{2}^{2}+N^{-\theta_{2}}\right)\,ds,$ (58)
+where $\theta_{2}\doteq\beta\gamma^{2}/d$ and $\gamma\in(0,1/2)$.
 2.2 Time evolution of energy
 Applying Itô’s lemma, following the approach of [25], we obtain
-$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}$ $\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}+\widetilde{I}^{1}_{t}-\widetilde{M}_{t}^{N}-2\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
+$\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}$ $\leq\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}+\widetilde{I}^{1}_{t}-\widetilde{M}_{t}^{N}-2\int_{0}^{t}\left\|\nabla(\rho_{s}^{N}-\rho_{s})\right\|_{2}^{2}\,ds$
 $+TN^{-\widetilde{\theta}_{1}}\,\|\nabla V\|_{2}^{2},$ (59)
@@
 $\widetilde{M}^{N}_{t}\doteq\frac{2\sqrt{2}}{N}\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{R}^{d}}(\rho_{s}-\rho_{s}^{N})\,(\nabla V^{N})(x-X_{s}^{i,N})\,dx\,dW_{s}^{i,N}.$
-3.2 Estimates for $\widetilde{I}_{t}$: the nonlinear terms
-Now by by subtracting and adding the term $\rho_{s}^{N}K\ast\rho_{s}^{N}$, along with integration by parts, we have
-$2\int_{\mathbb{R}^{2}}(\rho_{s}-\rho_{s}^{N})\left[\nabla\cdot\left[\rho_{s}K\ast\rho_{s}\right]-\nabla\cdot\left[\left\langle S_{s}^{N},V^{N}(x-\cdot)K\ast\rho_{s}^{N}(\cdot)\right\rangle\right]\right]\,dx$
-$=-2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left[\rho_{s}K\ast\rho_{s}-\left\langle S_{s}^{N},V^{N}(x-\cdot)K\ast\rho_{s}^{N}(\cdot)\right\rangle\right]\,dx$
-$=-2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left[\rho_{s}K\ast\rho_{s}-\rho_{s}^{N}K\ast\rho^{N}_{s}\right]\,dx$
+2.2.1 Estimates for $\widetilde{I}_{t}$: the nonlinearity
+We now address the nonlinearity in the evolution equation of energy.
+The main novelty is identifying, within the energy estimates, the Fisher information of the particle system alongside Ladyzhenskaya inequality, to address the nonlinearity.
+By subtracting and adding the term $\rho_{s}^{N}K\ast\rho_{s}^{N}$, we have
+$\widetilde{I}^{1}_{t}=-2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left[\rho_{s}K\ast\rho_{s}-\rho_{s}^{N}K\ast\rho^{N}_{s}\right]\,dx$
 $+2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})\left\langle S_{s}^{N},V^{N}(x-\cdot)\left[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right]\right\rangle\,dx$
 $\doteq\widetilde{I}^{1,1}+\widetilde{I}^{1,2}.$ (60)
-In addition, since $\nabla\cdot K=0$, by by subtracting and adding the terms $\rho_{s}$ and $\rho_{s}^{N}K\ast\rho_{s}$, we find
+In addition, since $\nabla\cdot K=0$, by subtracting and adding the term $\rho_{s}$, we find
 $\left|\widetilde{I}^{1,1}\right|$ $\leq\left|\underbrace{2\int_{\mathbb{R}^{2}}\nabla(\rho_{s}-\rho_{s}^{N})(\rho_{s}-\rho_{s}^{N})K\ast\rho_{s}\,dx}_{=0}\right|$
@@
 […]
-By Young’s inequality, we observe that
+By Young inequality, we observe that
 $\|\rho_{s}-\rho_{s}^{N}\|_{1}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}\leq C_{\epsilon}\|\rho_{s}\|_{2}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (62)
 By Ladyzhenskaya and Young inequalities, we deduce
-$\|\rho_{s}-\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$ $\lesssim\|\rho_{s}-\rho_{s}^{N}\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
-$\leq C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (88)
+$\|\rho_{s}-\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
+$\hskip 80.29999pt\lesssim\|\rho_{s}-\rho_{s}^{N}\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{1/2}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}\|\rho_{s}\|_{2}$
+$\hskip 80.29999pt\leq C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (63)
 Therefore by (62) and (63) in (61), we get
 $\widetilde{I}^{1,1}\leq C_{\epsilon}\|\rho_{s}\|_{2}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+2\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}.$ (64)
-Now we will take care about $\widetilde{I}^{1,2}$ in (85). First, taking into account (7), with analogous computations such as in section 2 of [40], we derive
+Now we will take care about $\widetilde{I}^{1,2}$ in (60).
+First, by decay of $V$, given by (6), we have, for $|x-\cdot|\geq N^{-\beta\gamma/d}$
+$V^{N}(x-\cdot)|x-\cdot|^{\gamma}$ $=N^{\beta}V(N^{\beta/d}(x-\cdot))|x-\cdot|^{\gamma}$
+[…]
+$\lesssim N^{-3+\beta}\left|(x-\cdot)\right|^{-3d/\beta}$
+$\lesssim N^{-\widetilde{\theta}_{2}},$ (65)
+which implies that
 $\left|\left\langle S^{N}_{s},V^{N}(x-\cdot)|x-\cdot|^{\gamma}\right\rangle\right|\leq C_{d}N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}(x)$ (66)
-with $\widetilde{\theta}_{2}\doteq\beta(2-3\gamma)$ and $\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{d}$, for $\gamma\in(1/2,2/3)$.
+with $\widetilde{\theta}_{2}\doteq 3-\beta-3\gamma$ and $\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{d}$, for $\gamma\in(0,1/2)$.
 It follows by (7), that
@@
 […]
-$\lesssim 2\|\rho_{s}^{N}\|_{1}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)$
-$+2\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)$
+$\lesssim 2\|\rho_{s}^{N}\|_{1}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)\,dx$
+$+2\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\left(N^{-\widetilde{\theta}_{2}}+N^{-\widetilde{\theta}_{3}}\rho_{s}^{N}\right)\,dx$
 $\doteq\widetilde{I}^{1,2,1}+\widetilde{I}^{1,2,2}$ (67)
-Now, by by subtracting and adding $\rho_{s}$, we note that
+Now, by subtracting and adding $\rho_{s}$ along side Young inequality, we note that
 $\widetilde{I}^{1,2,1}$ $\leq N^{-\widetilde{\theta}_{2}}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)$
-$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}-\rho_{s}^{N}|$
-$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}|$
+$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}-\rho_{s}^{N}|\,dx$
+$+N^{-\widetilde{\theta}_{3}}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})||\rho_{s}|\,dx$
 $\overset{\text{Young}}{\leq}N^{-\widetilde{\theta}_{2}}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)$
@@
 Again, by by subtracting and adding $\rho_{s}$, we find
-$\widetilde{I}^{1,2,2}$ $\leq N^{-\widetilde{\theta}_{2}}\|\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}+N^{-\widetilde{\theta}_{3}}\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\rho_{s}^{N}$
+$\widetilde{I}^{1,2,2}$ $\leq N^{-\widetilde{\theta}_{2}}\|\rho_{s}^{N}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}+N^{-\widetilde{\theta}_{3}}\|\rho_{s}^{N}\|_{4}\int_{\mathbb{R}^{2}}|\nabla(\rho_{s}-\rho_{s}^{N})|\rho_{s}^{N}\,dx$
 $\overset{\text{Holder}}{\leq}N^{-\widetilde{\theta}_{2}}\|\rho_{s}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}+N^{-\widetilde{\theta}_{2}}\|\rho_{s}^{N}-\rho_{s}\|_{4}\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{1}$
@@
 $+N^{-\widetilde{\theta}_{3}}\left(C_{\epsilon}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{6}+\epsilon\|\nabla(\rho_{s}-\rho_{s}^{N})\|_{2}^{2}\right).$ (74)
-3.3 End of the proof: application of Gronwall Lemma
-First, with the same computations as in [40], for all $\delta\in(0,\widetilde{\theta}_{4})$ there exists a random variable $\widetilde{A}_{0}$, finite almost surely, such that
+2.3 Combining entropy and energy
+First, with the same computations as in [25], for all $\delta\in(0,\theta_{3})$ there exists a random variable $A_{0}$, finite almost surely, such that
+$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{3}+\delta},$ (75)
+with $\theta_{3}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
+In addition, with the same computations as in [36], for all $\delta\in(0,\widetilde{\theta}_{4})$ there exists a random variable $\widetilde{A}_{0}$, finite almost surely, such that
 $\sup_{t\in[0,T]}|\widetilde{M}_{t}^{N}|\leq\widetilde{A}_{0}N^{-\widetilde{\theta}_{4}+\delta},$ (76)
 with $\widetilde{\theta}_{4}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
-Then by (98) and (99) in (84), for $\epsilon<<1$, we obtain
-$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$ $\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}$
-$+C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\,ds$
+[…]
+$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\leq C\int_{0}^{t\wedge\tau^{N}}\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho^{N}_{s}-\rho_{s}\|_{2}^{2}\right)\,ds$
+$\overset{}{+}(-1+2\epsilon)\int_{0}^{t\wedge\tau^{N}}\left(\mathcal{I}(\rho_{s}^{N}|\rho_{s})+\|\nabla(\rho_{s}^{N}-\rho_{s})\|_{2}^{2}\right)\,ds$
+$+CN^{-\theta_{1}}+CN^{-\theta_{2}}+A_{0}N^{-\theta_{3}+\delta},$ (77)
+for a universal constant $C>0$.
+Analogously, by (74) and (76) in (59), we obtain
+$\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}-\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}$ $\leq C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds$
 $+\int_{0}^{t}\left(C_{\epsilon}+\|\rho_{s}\|_{4}^{2}+\|\rho_{s}\|_{2}^{4}\right)\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\,ds$
 $+C_{\epsilon}\int_{0}^{t}\|\rho_{s}-\rho_{s}^{N}\|_{2}^{6}\,ds$
-$\overset{}{+}N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(\|\nabla\rho_{s}\|_{1}+\|\nabla\rho_{s}^{N}\|_{1}\right)\,ds$
-$\overset{}{+}2N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(\|\nabla\rho_{s}\|_{1}^{2}+\|\nabla\rho_{s}^{N}\|_{1}^{2}\right)\,ds$
-$+N^{-\widetilde{\theta}_{3}}C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\,ds+N^{-\widetilde{\theta}_{1}}\int_{0}^{t}\|\rho_{s}\|_{4}^{2}\,ds$
+$\overset{}{+}2N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(\|\nabla\rho_{s}\|_{1}^{2}+\|\nabla\rho_{s}^{N}\|_{1}^{2}+T\right)\,ds$
+$+N^{-\widetilde{\theta}_{3}}C_{\epsilon}\int_{0}^{t}\|\rho_{s}\|_{2}^{2}\,ds+N^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\rho_{s}\|_{4}^{2}\,ds$
 $+N^{-\widetilde{\theta}_{3}}\int_{0}^{t}\left(C_{\epsilon}\|\rho_{s}\|_{4}^{4}+C_{\epsilon}\|\rho_{s}\|_{2}^{4}\right)\,ds$
+$+(-2+8\epsilon)\int_{0}^{t}\left\|\nabla(\rho_{s}^{N}-\rho_{s})\right\|_{2}^{2}\,ds$
 $+TN^{-\widetilde{\theta}_{1}}\,\|\nabla V\|_{2}^{2}$
 $+\widetilde{A}_{0}N^{-\widetilde{\theta}_{4}+\delta}.$ (78)
-Now by (18) $\|\rho\|_{T,2},\|\rho\|_{T,4}\leq C_{0}$, almost surely, and thus for a universal constant $C>0$ in (100) along side (60), (2.3) and Csiszár-Kullback-Pinsker inequality (107), we derive
-$\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
-$\hskip 100.37498pt\leq\left\|\rho_{0}-\rho^{N}_{0}\right\|_{2}^{2}$
+Now, since $\rho_{s}^{N}\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, $\mathbb{P}$-a.s. and then by Holder’s inequality we obtain
+$\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\,dx$ $=\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\frac{\sqrt{\rho_{s}^{N}}}{\sqrt{\rho_{s}^{N}}}\,dx$
+$\overset{\text{Holder}}{\leq}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}^{d}}\rho_{s}^{N}\,dx\right)^{\frac{1}{2}}$
+$\overset{}{=}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}.$ (79)
+Also, from Leibniz’s rule we deduce,
+$\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx$ $=\int_{\mathbb{R}^{d}}\frac{\left|\nabla\left(\rho_{s}\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
+$\leq 2\int_{\mathbb{R}^{d}}\frac{\left|\nabla\rho_{s}\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx+2\int_{\mathbb{R}^{d}}\frac{\left|\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
+$\overset{}{\leq}2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\frac{\left|\nabla\rho_{s}\right|^{2}}{\rho_{s}^{2}}\,dx+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx$
+$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (80)
+where in the last step we are using the following: for $\eta>2C_{1}(8T+C_{3})+1$
+$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty$
+which implies by Donsker-Varadhan inequality (130),
+[…]
+[…]
+Hence since (17) implies $\|\rho\|_{L^{\infty}L^{2}},\|\rho\|_{L^{\infty}L^{4}}\leq C_{0}$, almost surely, by (78), (79) and (81), there exists a universal constant $C>0$ in (78), we derive
+$\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}^{N}-\rho_{s})\right\|_{2}^{2}\,ds$
+$\hskip 100.37498pt\leq\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}$
 $\hskip 100.37498pt+\int_{0}^{t}C\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\int_{0}^{t}C\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}\,ds$
 $\hskip 100.37498pt+\int_{0}^{t}C\|\rho_{s}-\rho_{s}^{N}\|_{2}^{6}\,ds$
-$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\nabla\rho_{s}\|_{1}\,ds+CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\nabla\rho_{s}\|_{1}^{2}\,ds$
-$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{1}}$
+$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\|\nabla\rho_{s}\|_{1}^{2}\,ds+TCN^{-\widetilde{\theta}_{2}}$
 $\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{2}}\int_{0}^{t}\left(2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+1+2\mathcal{I}(\rho_{s}^{N}|\rho_{s})\right)\,ds$
-$\hskip 100.37498pt+CN^{-\widetilde{\theta}_{3}}+CN^{-\widetilde{\theta}_{2}}$
+$\hskip 100.37498pt\overset{}{+}CN^{-\widetilde{\theta}_{1}}+CN^{-\widetilde{\theta}_{2}}+CN^{-\widetilde{\theta}_{3}}$
 $\hskip 100.37498pt+CN^{-\widetilde{\theta}_{1}}\,\|\nabla V\|_{2}^{2}$
 $\hskip 100.37498pt+C\widetilde{A}_{0}N^{-\widetilde{\theta}_{4}+\delta}.$ (82)
-Finally, by adding (65) and (101), we deduce for $\epsilon<<1$ and $N>\left(\frac{4}{\epsilon}\right)^{1/\widetilde{\theta}_{1}}$
+It follows that, by adding (77) and (82), for $\epsilon<<1$ and $N>>1$, we deduce
 $\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})+\left\|\rho_{t\wedge\tau^{N}}-\rho^{N}_{t\wedge\tau^{N}}\right\|_{2}^{2}+\int_{0}^{t\wedge\tau^{N}}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
-$\leq\mathcal{H}(\rho_{0}^{N}|\rho_{0})+\|\rho_{0}-\rho_{0}^{N}\|_{2}^{2}+C_{N}+TC$
+$\leq\mathcal{H}(\rho_{0}^{N}|\rho_{0})+\|\rho_{0}-\rho_{0}^{N}\|_{2}^{2}+C_{N}$
 $+\int_{0}^{t\wedge\tau^{N}}C\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\left(\int_{0}^{s}\left\|\nabla(\rho_{r}-\rho^{N}_{r})\right\|_{2}^{2}\,dr\right)\right)\,ds$
@@
 with
-$C_{N}\doteq CN^{-\widetilde{\theta}}\|\nabla\rho\|_{T,1}+CN^{-\widetilde{\theta}}\|\nabla\rho\|_{T,1}^{2}+CN^{-\widetilde{\theta}}+CA_{0}N^{-\widetilde{\theta}}+C\widetilde{A}_{0}N^{-\widetilde{\theta}}$,
+$C_{N}\doteq CN^{\widetilde{\theta}}\|\nabla\rho\|_{L^{\infty}L^{1}}^{2}+CN^{\widetilde{\theta}}+CA_{0}N^{\widetilde{\theta}}+C\widetilde{A}_{0}N^{\widetilde{\theta}}$,
 and
-$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
+$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\theta_{3},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
 for $\delta<<1$.
-It follows by (15), (23), (24) and (81), (99) in (102), that there exist $T>0$ and $\widetilde{\Lambda}\in\mathcal{F}$, $\mathbb{P}(\widetilde{\Lambda})=1$, with the following property: for all $\omega\in\widetilde{\Lambda}$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have $N^{-\widetilde{\theta}}+C_{N}+TC<\exp{(-CT)}(2CT)^{-1/2}$ and
+2.4 End of the proof: Removal the stopping time
+We now remove the stopping time $\tau^{N}$ introduced in (32) by exploiting the probabilistic framework provided by Assumptions $\left(\mathbf{A}^{\rho_{0}}\right)$ and $\left(\mathbf{A}^{X_{0}}\right)$.
+It allows us to pass from the localized estimates to uniform bounds with high probability, removing the stopping time.
+First, since $K=\nabla\cdot K_{0}$, $K_{0}\in L^{\infty}$ and $V^{N}=N^{\beta}V(N^{\beta/d}\cdot)$, by (2), we have
+$\sup_{t\in[0,T]}|X_{t}^{i,N}|$ $\leq|X_{0}^{i,N}|+T\|\nabla V\|_{1}\|K_{0}\|_{\infty}N^{\beta}$
+$+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sqrt{2}dW_{s}^{i,N}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|,$ (84)
+which implies for $\widetilde{C}_{K_{0},V}\doteq(1-T\|\nabla V\|_{1}\|K_{0}\|_{\infty})$
+$\{\tau^{N}<T\}$ $\subset\bigcup_{i=1}^{N}\left\{\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
+$=\left\{\exists i;\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
+$\subset\left\{\exists i;|X_{0}^{i,N}|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sqrt{2}dW_{s}^{i,N}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|\geq\widetilde{C}_{K_{0},V}N^{\beta}\right\}$
+$\subset L_{1}^{N}\cup L_{2}^{N}\cup L_{3}^{N},$ (85)
+with
+$L_{1}^{N}\doteq\left\{\exists i;|X_{0}^{i,N}|\geq C_{K_{0},V}N^{\beta}\right\},$ (86)
+$L_{2}^{N}\doteq\left\{\exists i;\sup_{t\in[0,T]}\left|\int_{0}^{t}dW_{s}^{i,N}\right|\geq C_{K_{0},V}N^{\beta}\right\}$ (87)
+and
+$L_{3}^{N}\doteq\left\{\exists i;\sup_{t\in[0,T]}\left|\int_{0}^{t}\sigma_{s}dB_{s}\right|\geq C_{K_{0},V}N^{\beta}\right\}$ (88)
+with $C_{K_{0},V}\doteq(1-T\|\nabla V\|_{1}\|K_{0}\|_{\infty})/3\sqrt{2}>0$, for $T<(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})$.
+We note that, with the same computations like those led (34), we deduce
+$\mathbb{P}\left(L_{1}^{N}\right)$ $\overset{}{\leq}N(C_{K_{0},V}N^{\beta})^{d-2}e^{-(C_{K_{0},V}N^{\beta})^{2}}$ (89)
+We now take care about the $L_{2}^{N}$. First, we observe that, in view of Definition 5.1 in Chapter 2 of [24],
+$\Omega=\prod_{j=1}^{d}\Omega^{j},\,\,\,\mathcal{F}=\bigotimes_{j=1}^{d}\mathcal{F}^{j},\,\,\,\mathbb{P}=\bigotimes_{j=1}^{d}\mathbb{P}^{j}$ (90)
+and then
+$W_{t}^{i,N}(\omega)=(W_{t}^{i,N,j}(\omega_{j}))_{j=1}^{d}$ (91)
+for $W^{i,N,j}$ a standard, one-dimensional Brownian motion on $(\Omega^{j},\mathcal{F}^{j},\mathbb{P}^{j})$, $j\in\{1,...,d\}$.
+Therefore, we note that $W_{t}^{i,N}=\int_{0}^{t}dW_{s}^{i,N}$, and since
+$|W_{t}^{i,N}|^{2}=\sum_{j=1}^{d}(W_{t}^{i,N,j})^{2},$
+if $\sup_{t\leq T}|W_{t}^{i,N}|\geq a$, then there exists $j\in\{1,...,d\}$, such that $\sup_{t\leq T}|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}.$
+Hence,
+$\left\{\sup_{t\leq T}|W_{t}^{i,N}|\geq a\right\}\subset\bigcup_{j=1}^{d}\left[\Omega^{1}\times...\times\widetilde{\Omega}^{j}\times...\times\Omega^{d}\right]$ (92)
+with
+$\widetilde{\Omega}^{j}\doteq\left\{\sup_{t\leq T}|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}\right\}$
+and then, we obtain
+[…]
+$\leq\left(\sqrt{\frac{T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}\right)\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2T)\right)},$ (93)
+where in the second inequality, we are using (8.3)’ in Chapter 2 of [24], page 96.
+Now we will derive an analogous estimate for $L_{3}^{N}$. First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$. Define,
+$Y_{t}\doteq\int_{0}^{t}\sigma_{s}\,dB_{s},$ (94)
+which implies
+$Y_{t}^{j}=\sum_{k=1}^{d}\underbrace{\int_{0}^{t}\sigma_{s}^{j,k}\,dB_{s}^{k}}_{\doteq Y_{s}^{j,k}}.$ (95)
+Since $\sigma:[0,T]\to\mathbb{R}^{d}\times\mathbb{R}^{d}$ is bounded, for $j\in\{1,...,d\}$, the function defined by $t\mapsto\langle Y^{j,k}\rangle_{t}$ is continuous and increasing.
+Hence, by Theorem 4.6, in Chapter 3 of [24], for all $j\in\{1,...,d\}$, we get
+$\sup_{t\leq T}|Y_{t}^{j}|$ $\leq\sum_{k=1}^{d}\sup_{t\leq T}\left|\int_{0}^{t}\sigma_{s}^{j,k}\,dB_{s}^{k}\right|\leq\sum_{k=1}^{d}\sup_{t\leq T}\left|\widetilde{B}^{j,k}_{\langle Y^{j,k}\rangle_{t}}\right|$
+$\leq\sum_{k=1}^{d}\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j,k}_{t}|,$ (96)
+where $(\widetilde{B}^{j,k}_{t})_{t\geq 0}$ is the time-change process, which in particular, is a one dimensional standard Brownian motion. In addition, since
+$|Y_{t}|^{2}=\sum_{j=1}^{d}|Y_{t}^{j}|^{2}$
+if $\sup_{t\leq T}|Y_{t}|\geq a$, then there exists $j\in\{1,...,d\}$, such that $\sup_{t\leq T}|Y_{t}^{j}|\geq\frac{a}{\sqrt{d}}.$ Therefore, by (90) and for all $a>0$, we find
+$\left\{\sup_{t\leq T}|Y_{t}|\geq a\right\}$ $\subset\bigcup_{j=1}^{d}\left\{\sup_{t\leq T}|Y_{t}^{j}|\geq\frac{a}{\sqrt{d}}\right\}$
+[…]
+$\overset{}{\subset}\bigcup_{j=1}^{d}\bigcup_{k=1}^{d}\left[\Omega^{1}\times...\times\widetilde{\Omega}^{j,k}\times...\times\Omega^{d}\right],$ (97)
+with
+$\widetilde{\Omega}^{j,k}\doteq\left\{\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j,k}_{t}|\geq\frac{a}{d^{3/2}}\right\}.$ (98)
+It follows that, by (96) and (97)
+[…]
+$\leq\left(\sqrt{\frac{\|\sigma\|_{\infty}^{2}T}{2\pi}}\frac{4d^{2}N}{\frac{C_{K_{0},V}N^{\beta}}{d^{3/2}}}\right)\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{d^{3/2}}\right)^{2}/(2\|\sigma\|_{\infty}^{2}T)\right)},$ (99)
+where in the second inequality, we are using (8.3)’ in Chapter 2 of [24], page 96.
+Hence, by (89), (93) and (99) in (85), we find
+$\sum_{N\in\mathbb{N}}\mathbb{P}\left(\tau^{N}<T\right)<\infty$ (100)
+and then by Borel-Cantelli Lemma, we arrive at
+$\mathbb{P}\left(\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}\{\tau^{N}<T\}\right)=0.$ (101)
+Finally, by (14), (23), (19), (20), (37), (75), (76) and (101) in (83) , for $T<\min{\left(T_{1};(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})\right)}$ for all $\omega\in\Lambda$, with
+$\Lambda\doteq\Lambda_{0}\cap\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}\{\tau^{N}=T\}\right)\cap\left\{A_{0}<\infty\right\},$ (102)
+there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, we have $\tau^{N}(\omega)=T$,
+$\mathcal{H}(\rho_{0}^{N}|\rho_{0})+\|\rho_{0}-\rho_{0}^{N}\|_{2}^{2}+C_{N}<\exp{(-CT)}(2CT)^{-1/2}$
+and
 $\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
-$\hskip 20.075pt\leq N^{-\widetilde{\theta}}+C_{N}+TC$
+$\hskip 20.075pt\leq\mathcal{H}(\rho_{0}^{N}|\rho_{0})+\|\rho_{0}-\rho_{0}^{N}\|_{2}^{2}+C_{N}$
 $\hskip 20.075pt+\int_{0}^{t}C\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\left(\int_{0}^{s}\left\|\nabla(\rho_{r}-\rho^{N}_{r})\right\|_{2}^{2}\,dr\right)\right)\,ds$
 $\hskip 20.075pt+\int_{0}^{t}C\left(\mathcal{H}(\rho_{s}^{N}|\rho_{s})+\|\rho_{s}-\rho_{s}^{N}\|_{2}^{2}+\left(\int_{0}^{s}\left\|\nabla(\rho_{r}-\rho^{N}_{r})\right\|_{2}^{2}\,dr\right)\right)^{3}\,ds.$ (103)
-In particular by the nonlinear Gronwall’s inequality (111), for all $t\in[0,T]$, we deduce
-$\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$ $\leq N^{-\widetilde{\theta}}+C_{N}+TC$ (104)
+In particular by the nonlinear Gronwall inequality (133) in (103), for all $t\in[0,T]$, we deduce
+$\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{t}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
+$\hskip 80.29999pt\leq\mathcal{H}(\rho_{0}^{N}|\rho_{0})+\|\rho_{0}-\rho_{0}^{N}\|_{2}^{2}+C_{N}.$ (104)
 and thus, we end up with
-$\limsup_{N\to\infty}\sup_{t\in[0,T]}\left(\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}-\rho^{N}_{t}\right\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds\right)$ $\lesssim T,$ (105)
+$\lim_{N\to\infty}N^{\widetilde{\theta}}\left(\sup_{t\in[0,T]}\left(\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}\right)+\int_{0}^{T}\left\|\nabla(\rho_{s}^{N}-\rho_{s})\right\|_{2}^{2}\,ds\right)$ $=0,$ (105)
 almost surely.
+3 Proof of Theorem 2
+In this section we present the proof of Theorem 2. In view of the previous section we only need the address the nonlinearity, for $d\geq 2$.
+3.1 Estimates for $I_{t}$: the nonlinearity
+We now address the nonlinear terms in (26).
+The main novelty is to use Donsker-Varadhan inequality (130), combined with the notion of solution of limiting process given by Definition 1.1, in the context of moderate interactions, along with the dissipation that comes from of Fisher information and Assumption $(\mathbf{A}^{\nabla\cdot K})$, to tackle with the singularity of kernel $K$.
+First, we set
+$I_{t}$ $\overset{}{=}\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho_{s}^{N}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)[K\ast\rho_{s}-K\ast\rho_{s}^{N}]\,dx\,ds$
+$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\underbrace{\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle}_{\doteq I^{N}}\,dx\,ds$
+$\doteq I^{1}_{t}-I^{2}_{t},$ (106)
+where
+$I^{1}_{t}$ $=-\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\rho_{s}^{N}\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
+$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\nabla\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
+$\doteq-I^{1,1}_{t}-I^{1,2}_{t}.$ (107)
+Now by $\epsilon$-Young inequality and convolution inequality, we deduce
+$I_{t}^{1,2}=\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\nabla\rho_{s}\left(\frac{\sqrt{\rho_{s}^{N}}}{\rho_{s}}\right)\left(\frac{\rho_{s}}{\sqrt{\rho_{s}^{N}}}\right)\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right][K_{0}\ast\rho_{s}-K_{0}\ast\rho_{s}^{N}]\,dx\,ds$
+$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\left|K_{0}\ast(\rho_{s}-\rho_{s}^{N})\right|^{2}\,dx\,ds$
+$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}dx\,ds+C_{\epsilon,K_{0}}\int_{0}^{t}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}dx\,ds.$ (108)
+We observe that, for $\eta>2C_{1}(8T+C_{3})+1$
+$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty$ (109)
+which implies by Donsker-Varadhan inequality (130),
+[…]
+[…]
+Hence by Csiszár-Kullback-Pinsker inequality (129) and (108), we derive
+$I_{t}^{1,2}$ $\leq\epsilon\int_{0}^{t}\mathcal{I}(\rho^{N}_{s}|\rho_{s})\,ds+C_{\epsilon,\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds.$ (111)
+Regarding term $I_{t}^{1,1}$ in (107), by convolution inequality, we find
+$I_{t}^{1,1}$ $\leq\int_{0}^{t}\|K_{0}\|_{\infty}\|\rho^{N}_{s}-\rho_{s}\|_{1}\int_{\mathbb{R}^{d}}\left|\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right|\rho^{N}_{s}\,dx\,ds.$ (112)
+In addition, if $C_{1,2}\doteq\max{(C_{1},C_{2})}$ since
+[…]
+for $\eta>2C_{1,2}(8T+C_{3})+1$, we obtain
+$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty.$ (114)
+It follows by Donsker-Varadhan inequality (130),
+[…]
+[…]
+Therefore, by (115) in (112), we get
+$I_{t}^{1,1}$ $\overset{}{\leq}C_{\eta,K_{0}}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (116)
+[…]
+$I_{t}^{1}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+\epsilon\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (117)
+Now, we will focus on the term $I_{t}^{2}$ in (44). First, we recall that by Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$, $K=\nabla\cdot K_{0}$ with $K_{0}\in L^{\infty}$. Hence, we arrive at
+$\left|K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right|$ $=\left|K_{0}\ast\nabla\rho_{s}^{N}(\cdot)-K_{0}\ast\nabla\rho_{s}^{N}(x)\right|$
+$\leq\int_{\mathbb{R}^{d}}\left|K_{0}(\cdot-y)-K_{0}(x-y)\right||\nabla\rho_{s}^{N}(y)|\,dy$
+$\leq\int_{\mathbb{R}^{d}}2\|K_{0}\|_{\infty}|\nabla\rho_{s}^{N}(y)|\,dy$
+$=2\|K_{0}\|_{\infty}\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}(y)|\,dy.$ (118)
+Additionally, $\rho_{s}^{N}\in\mathcal{P}\left(\mathbb{R}^{d}\right)$, $\mathbb{P}$-a.s. and then by Holder’s inequality we obtain
+$\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\,dx$ $=\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}|\frac{\sqrt{\rho_{s}^{N}}}{\sqrt{\rho_{s}^{N}}}\,dx$
+$\overset{\text{Holder}}{\leq}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}^{d}}\rho_{s}^{N}\,dx\right)^{\frac{1}{2}}$
+$\overset{}{=}\left(\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx\right)^{\frac{1}{2}}.$ (119)
+Now from Leibniz’s rule we deduce,
+$\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}^{N}|^{2}}{\rho_{s}^{N}}\,dx$ $=\int_{\mathbb{R}^{d}}\frac{\left|\nabla\left(\rho_{s}\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
+$\leq 2\int_{\mathbb{R}^{d}}\frac{\left|\nabla\rho_{s}\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx+2\int_{\mathbb{R}^{d}}\frac{\left|\rho_{s}\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}}{\rho_{s}^{N}}\,dx$
+$\overset{}{\leq}2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\frac{\left|\nabla\rho_{s}\right|^{2}}{\rho_{s}^{2}}\,dx+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx$
+[…]
+[…]
+$I^{N}$ $\leq 2\|K_{0}\|_{\infty}\left\langle S_{s}^{N},V^{N}(x-\cdot)\right\rangle\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}$
+$\leq 1/2\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}},$
+and then
+$I^{2}_{t}$ $\doteq\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
+$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
+$\leq\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})\right)^{\frac{1}{2}}\,dx\,ds$
+$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)\rho^{N}_{s}\left(2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\,dx\,ds$
+$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds.$ (121)
+It follows that, by Young and Holder inequalities,
+$I_{t}^{2}\overset{\text{Young}}{\leq}\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds$
+$+\sqrt{2}/2\int_{0}^{t}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\,dx\,ds$
+$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\epsilon,\eta}$
+$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\epsilon,\eta}.$ (122)
+[…]
+$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\epsilon,\eta}.$ (123)
+3.2 Application of Gronwall’s Lemma
+[…]
+$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds$
+$\overset{}{+}(3\epsilon+\sqrt{2}/2)\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
+$-\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
+$+C_{\alpha,T}N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+TC_{\eta}$ (124)
+and then for $\eta>>1$ and $\epsilon<<1$, we deduce
+$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+T,$ (125)
+with $\theta_{1}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$ and $\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
+Finally, from Gronwall’s inequality (131) in (125), with the analogous computations as led (102), there exists $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the property that, for all $\omega\in\Lambda$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)=T$ and
+$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right)$ (126)
+with
+$\theta\doteq\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta\right).$
+In particular, we get
+$\limsup_{N\to\infty}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim T,\,\,\,\,\,\,\mathbb{P}-a.s.$ (127)
 4 Appendix